### v1

# SoK: LLM-based Log Parsing

Abstract.
Log data, generated by software systems, provides crucial insights for tasks like monitoring, root cause analysis, and anomaly detection. Due to the vast volume of logs, automated log parsing is essential to transform semi-structured log messages into structured representations. Traditional log parsing techniques often require manual configurations, such as defining log formats or labeling data, which limits scalability and usability. Recent advances in large language models (LLMs) have introduced the new research field of LLM-based log parsing, offering potential improvements in automation and adaptability. Despite promising results, there is no structured overview of these approaches since this is a relatively new research field with the earliest advances published in late 2023. This paper systematically reviews $29$ LLM-based log parsing methods, comparing their capabilities, limitations, and reliance on manual effort. We analyze the learning and prompt-engineering paradigms employed, efficiency- and effectiveness-enhancing techniques, and the role of LLMs in the parsing process. We aggregate the results of the survey in a large table comprising the characterizing features of LLM-based log parsing approaches and derive the general process of LLM-based log parsing, incorporating all reviewed approaches in a single flow chart. Additionally, we benchmark seven open-source LLM-based log parsers on public datasets and critically assess their reproducibility. Our findings summarize the advances of this new research field and provide insights for researchers and practitioners seeking efficient and user-friendly log parsing solutions, with all code and results made publicly available for transparency.
parsing, log data, large language models
copyright: nonedoi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xpublicationmonth: X
1. Introduction
Log data consists of recorded information generated by logging statements in software applications, systems, devices, or networks during operation. It provides a chronological record of events and contains valuable insights for various tasks, such as monitoring (Chen et al., 2019; Wang et al., 2021), root cause analysis (He et al., 2021; Notaro et al., 2023), and anomaly detection (Landauer et al., 2023a; He et al., 2016a). Given that modern computer systems can generate millions of log lines per hour, manual processing is infeasible, necessitating fast and scalable automated approaches. For instance, in 2013 the Alibaba Cloud system was reported to produce approximately 100 to 200 million log lines per hour (Mi et al., 2013). However, extracting meaningful information from raw log data requires transforming it into a structured format. Since logs are typically semi-structured, following general patterns but containing variable log messages, direct analysis is challenging. To enable log analytics tasks like anomaly detection, logs must first be parsed into a structured representation, making log parsing a critical step in log analysis (Wurzenberger et al., 2019).
Log parsing refers to the extraction of structured data from semi-structured logs. Parsing the logs into the structured format can simply be done by regular expression matching if the correct log templates are given. Most methods therefore formulate log parsing as a log template extraction problem (He et al., 2017; Jiang et al., 2024a). Ground-truth templates are defined somewhere in logging statements in the source code and are therefore often not available. A wide range of log parsing techniques (Shima, 2016; Jiang et al., 2008; Du and Li, 2016; Sedki et al., 2022; Le and Zhang, 2023b) have therefore been proposed to overcome this issue.
Many state of the art log parsing methods require manual configuration or other manual actions at some point in the parsing process. Approaches may involve labeling of some of the logs from the training set (Jiang et al., 2024a; Xu et al., 2024; Le and Zhang, 2023b), definition of log formats (Li et al., 2023) or regular expressions for preprocessing (Yu et al., 2023; He et al., 2017; Sedki et al., 2022), or other parameters (Yu et al., 2023) that fit the parsing algorithm to the data. From a user’s perspective the configuration of parsing algorithms might seem overwhelming since it requires expertise or an analysis of the data at hand. The reason log parsing methods require manual actions is that humans have both semantic and syntactic understanding (Jiang et al., 2024a), but more importantly, a natural generalization capability and potentially expert knowledge about logs and log parsing. Many log parsers have already achieved semantic (Le and Zhang, 2023b; Liu et al., 2022; Jiang et al., 2024a) and syntactic abilities (He et al., 2017; Dai et al., 2022; Yu et al., 2023). While syntax-based parsers focus on characteristics like occurrence frequencies, or word or log length, to determine the constant part of logs, semantic-based parsers leverage the semantic differences between static and variable parts of log messages (Jiang et al., 2024a; Xiao et al., 2024). Knowledge about logs and how to parse them can be learned by language models, such as BERT (Devlin et al., 2019) or its descendants, by fine-tuning them with logs and their templates (Le and Zhang, 2023b; Tao et al., 2022). However, these methods often lack generalization and have to be trained or fine-tuned with labeled logs, preferably with logs of the exact same log type as the target logs (Ma et al., 2024a). Generalization arises from pre-training on a large corpus of diverse datasets, allowing models to perform well also on novel tasks and possibly enhanced through fine-tuning (FT) (Gao et al., 2021) or in-context learning (ICL) (Brown et al., 2020). This is where large language models (LLMs) come into play.
With the emergence of LLMs the new research field of LLM-based log parsing arised. In late 2023, Le and Zhang (Le and Zhang, 2023a) reported notable performance of ChatGPT [^1] https://openai.com/index/chatgpt/; accessed 10-March-2025 in a naive setting, where the LLM was presented individual logs and simply asked for their templates. In the meantime, other approaches adopted LLMs for this purpose as well, building sophisticated frameworks that utilize LLMs in various ways to either parse logs directly or enhance log parsing, by adopting learning paradigms for LLMs such as ICL (Brown et al., 2020) or fine-tuning (Gao et al., 2021). Given the fact that the number of unique templates is significantly lower than the number of logs and the average rate of newly discovered templates decreases as more logs are processed (Jiang et al., 2024a), some LLM-based parsers (Jiang et al., 2024a) are even able to achieve runtime performance comparable with the fast and well-known parser Drain (He et al., 2017).
To the best of our knowledge, there is an absence of a structured overview and an objective comparison of papers and methods concerning LLM-based log parsing, as this is a rather novel research field which only recently emerged from the popular research field of generative LLMs. So far there is merely a survey paper about LLM-based log analysis techniques by Akhtar et al. (Akhtar et al., 2025), which provides a rudimentary overview on using LLMs for log parsing and the techniques employed. Consequently, this work aims to undertake a systematic review of the literature to identify the commonalities and variabilities of the existing approaches and to present our gained insights. Our main focus is an overview of LLM-based log parsing with emphasis on the users’ perspective — thus, how the LLM is applied and how much manual effort the approaches require. The aim is therefore to provide a comprehensive overview of the existing approaches and techniques, enabling anyone interested in (LLM-based) log parsing to quickly and easily identify the most suitable solution for their specific use case.
The aforementioned goals are formulated as the following research questions:
  • RQ1: What are the main advantages and disadvantages of LLM-based log parsing approaches and non-LLM-based approaches?
  • RQ2: To what extent do LLM-based log parsing approaches rely on labeled data and manual configuration?
  • RQ3: Which techniques can enhance the efficiency or effectiveness of LLM-based log parsing?
  • RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
  • RQ5: To what extent are the presented LLM-based parsing methods accessible and reproducible?
We summarize the main contributions of the present work as follows:
  • We provide a systematic overview of the existing methods, based on a set of feature definitions that characterize LLM-based log parsing approaches, or the corresponding papers, respectively.
  • We derive the general process of LLM-based log parsing, encompassing each parsing approach of each reviewed paper in a single flow chart.
  • We provide a comprehensive benchmark of seven open-source LLM-based log parsing approaches on open-source datasets.
  • Based on the literature review and our evaluation results, we answer $5$ research questions and derive further findings that may benefit the future research and design of log parsers.
  • Finally, we make the source code and all the results of our evaluation publicly available in our GitHub repository [^2] https://github.com/ait-aecid/LLM-log-parsing.git to ensure transparency and reproducibility.
The remainder is structured as follows: Section 2 describes the background and the core concepts and how we understand them. Section 3 outlines the survey method, including the search strategy and the definition of the reviewed features. Section 4 presents the results of the literature review in a large table and describes the findings of each feature. Section 5 explains the evaluation setup, while Sec. 6 presents the evaluation results. We discuss the findings from literature and evaluation in Sec. 7 and conclude our paper in Sec. 8.
2. Background and Related Work
In this Section, we describe the relevant background by means of the core concepts relevant to LLM-based log parsing.
2.1. Log Parsing
Log parsing is the foremost step in log analytics and denotes the process of extracting structured data from logs. Logs are given as single- or multi-line events and are available in textual form (Landauer et al., 2023a). The diverse and often un- or semi-structured nature of logs makes log parsing a difficult task. It is the precondition for many log analytics tasks, hence the effectiveness of downstream tasks is therefore inherently dependent on the effectiveness of the parser and also determines the runtime efficiency of the whole process, which is especially important for online (real-time) or incremental approaches, such as live-monitoring for cybersecurity (Jiang et al., 2024a). Incremental approaches are designed to handle continuous data streams, updating the parsing model dynamically as new log data arrives. This allows for real-time analysis and quicker responses to emerging issues.
In this work, we specifically focus on template extraction, since the extraction of the variables from the log messages is trivial once the (correct) templates are identified, for example, with regular expression matching. The template extraction process is therefore the key component in log parsing.
2.2. Log Templates
A log template is a predefined format used to log system events in software applications. It typically includes placeholders for essential details like timestamps, log levels (e.g., INFO, ERROR), message content, contextual metadata and more (Li et al., 2023). Previous research differs between the unstructured log message part of a log and the structured log headers (He et al., 2016b). The structured part is thereby straightforward to extract with regular expressions and it has been shown that parsing only the log message part improves the parsing accuracy (He et al., 2016b). Many methods (He et al., 2017; Du and Li, 2016; Xu et al., 2024; Du and Li, 2016), therefore require the log format of a dataset as input parameters. An example of a log template from the Apache dataset (Zhu et al., 2023) is given in Fig. 1. The log format parameter for this example would be “[<Time>] [<Level>] <Content>”, whereas “Time” and “Level” refers to log headers and “Content” refers to the log message. In the remainder, we understand “log format” as this parameter defining the positions of the log headers and of the content within different log types, or different datasets, respectively.
Figure 1. A simple example of log parsing, from logging statement to log file to parsed log.
2.3. Large Language Models (LLMs)
Large Language Models (LLMs) have evolved significantly, transitioning from early statistical and rule-based approaches to deep learning architectures powered by Generative Pretrained Transformers (GPTs). These models leverage self-attention mechanisms and large-scale datasets to generate human-like text, enabling advancements in natural language processing (NLP) tasks such as summarization, translation, content generation, and complex problem-solving. A key enhancement in LLMs is instruction tuning which specializes the pretrained LLMs to follow instructions provided in prompts. Instruction tuning enables effective ICL. Instead of modifying model parameters, ICL allows LLMs to learn from prompts that include instructions, examples, and queries, making them more adaptable to specialized applications (Zhang et al., 2024b).
The term LLM is only loosely defined, hence, for this work, we define LLMs as generative pretrained transformers (GPTs) such as ChatGPT, Mistral or Qwen. Models like BERT (Devlin et al., 2019) (encoder-only) are thereby excluded, as we are only interested in typically larger, generative models that excel in both understanding and generating text, making them suitable for a broader range of tasks. Many works included BERT into their definition of LLMs which complicated the manual filtering process after the keyword search.
2.4. Log Parsing with LLMs
Large-scale model series like ChatGPT and many more have demonstrated potential in this domain. Some approaches fine-tune smaller LLMs (Mehrabi et al., 2024) or BERT (Le and Zhang, 2023b) to this specific task while reporting superior parsing performance to state of the art approaches. However, fine-tuning LLMs for log parsing is resource-intensive in terms of computational costs, runtime, as well as labeled examples, making ICL a compelling alternative (Xiao et al., 2024). Research shows that ICL enhances LLMs’ performance in logical reasoning and fact retrieval, suggesting its viability for structured log analysis (Brown et al., 2020).
Despite these advantages, integrating LLMs into log parsing presents challenges. These models are not inherently designed for log parsing, leading to inconsistencies in token classification and errors in log templates. Such inaccuracies can negatively impact downstream tasks like anomaly detection (Zhu et al., 2019). Additionally, given the vast volume of real-world log data, efficient processing is crucial. The computational demands of LLMs contribute to inference overhead and network latency, necessitating optimizations for practical deployment (Jiang et al., 2024a).
3. Survey Method
3.1. Search Strategy
Subsequently, this section describes our search strategy and the inclusion and exclusion criteria derived from the core concepts of Sec. 2 and early insights from LLM-based log parsing approaches. For the keyword search, we determined the following keywords:
  1. (1)
    log(s) (in title);
  2. (2)
    LLM(s), large-language-model(s), or large language model(s);
  3. (3)
    parsing, parser(s), or parse.
The term ”log” or ”logs” must be included in the publication title, while the remaining keywords are entered into the default search engine of the respective database. This approach should ensure that the search results are focused on publications where log data is the central aspect. The search statistics are given in Table 1.
Table 1. Keyword search results (status 29-01-2025). “#R” and “#I” stand for the number of results and the number of papers that were still included after all exclusion criteria were applied.
Database Search String #R #I
IEEE Xplore ("Document Title": "log" OR "Document Title": "logs") AND (LLM OR LLMs OR "large language model" OR "large language models") AND (parser OR parsers OR parsing OR parse) 23 11
Science-Direct title: ("log?") AND title, abstract, keywords: ("LLM" OR "LLMs" OR "large language model" OR "large language models" OR "large-language-model" OR "large-language-models") AND ("parser" OR "parsers" OR "parsing" OR "parse") 6 0
ACM Digital Library [[Title: log] OR [Title: logs]] AND [[All: "llm"] OR [All: "llms"] OR [All: "large language model"] OR [All: "large language models"] OR [All: "large-language-model"] OR [All: "large-language-models"]] AND [[All: "parser"] OR [All: "parsers"] OR [All: "parsing"] OR [All: "parse"]] 40 9
Google Scholar intitle:"log" OR intitle:"logs" "LLM" OR "LLMs" OR "large-language-model" OR "large-language-models" OR "large language model" OR "large language models" "parsing" OR "parser" OR "parsers" OR "parse" 166 29
Web of Science TI=("log" OR "logs") AND TS=("LLM" OR "LLMs" OR "large language model" OR "large language models" OR "large-language-model" OR "large-language-models") AND TS=("parser" OR "parsers" OR "parsing" OR "parse") 2 2
Springer Link Title:(log OR logs) AND (LLM OR LLMs OR "large language model" OR "large language models" OR "large-language-model" OR "large-language-models") AND (parser OR parsers OR parsing OR parse) 6 0
The numbers of search results per database sum up to $176$, excluding duplicates found on multiple databases. The results were then filtered by exclusion criteria. A publication is excluded if,
  • It is not written in English.
  • It is not accessible in electronic format.
  • It is a book, technical report, lecture note, presentation, patent, thesis, or dissertation.
  • A more recent publication is available that presents the same study (by the same authors), ensuring that this SoK focuses on the latest versions of specific approaches.
  • It only applies an existing approach without introducing any novelties to LLM-based log parsing.
  • It does not apply log parsing or LLMs by definitions of Sec. 2.
  • It is not accessible with our institutions’ licenses.
By applying these exclusion criteria on both abstract and full text, we omitted $147$ of the initial publications found: $1$ publication was omitted because it was not written in English. $8$ publications were omitted due to being completely unrelated to the topic. $110$ are related but excluded, as they do not directly cover LLM-supported log data parsing by the definitions of Sec. 2 or are about other log analytics tasks, but employ parsing-free methods or existing conventional parsers, such as Drain (He et al., 2017) or SPELL (Du and Li, 2016). $21$ were omitted because they were books, theses, dissertations, or technical reports. $2$ were excluded for not being available freely with any license from our institutions. $5$ were outdated versions of newer papers.
The final selection consists of $29$ papers. One publication is from the year 2023, one from 2025, while the remaining are all from 2024. Since there is a recent hype around the usage of LLMs for log related issues, we include pre-print papers to capture the newest research. There are $10$ preprint papers in the final selection ($\sim 34\%$), all from $2024$.
3.2. Feature Definition
This section describes each feature and its classes of Table 2 which we report from the analyzed methods. We came up with this specific set of features, since they map to our research questions and pop up significantly often in many of the analyzed works. Additionally, we think that these are the most relevant for anyone who wants to apply LLM-based log parsing. The features either describe general properties (GP), processing steps (PS), or concern the reproducibility (R) of the approaches. For reproducibility, we take inspiration from the work of Olszewski et al. (Olszewski et al., 2023) where they analyzed machine learning papers of Tier 1 security conferences for reproducibility. The researchers defined a set of questions to determine the reproducibility of research findings.
In Table 2 features are reported with “✓” (true), an empty cell (false) or a descriptive name or abbreviation. If a paper provides an unclear answer to the feature in question, we write “?”. For instance, one papers provides a link to its code repository but the repository is empty, thus we write “?” for the corresponding feature. If a paper does not include the feature, or if the feature is not applicable, we leave it blank.
GP-1 — Supervision
In log parsing a labeled log is one for which a template, acting as the ground truth for that log, is available. We differentiate approaches based on the requirement for labels:
  • Supervised parsing (sup) requires at least some log instances of the training set to be labeled.
  • Unsupervised parsing (un) does not require labels.
GP-2 — Parsing Mode
Many methods are described as online approaches, but their interpretations vary. Some consider online processing to be incremental processing or streaming, while others apply batch-wise processing yet still label their method as online. Additionally, some works do not specify whether their approach is online or offline. Our initial incentive was to classify these methods accordingly, but this is often not feasible due to a lack of context or sufficient explanations. As a consequence, we devised the following categories that describe how many logs are processed at once:
  • Stream (S): The logs are processed one-by-one.
  • Batch (B): Multiple logs are processed at once, whereby it is possible to apply local (within each batch) optimizations. A batch is significantly smaller than the whole dataset.
  • Total (T): The entire dataset is processed at once, whereby it is possible to apply global optimizations to the process.
GP-3 — Learning / Prompting
We report four different types of learning or prompt engineering techniques. The type of learning can strongly influence how the prompt is designed which is why we report this in a single feature:
  • In-context learning (ZS/FS/S): ICL (Brown et al., 2020) leverages the context provided within a single LLM call to generate responses, adapting to specific needs without any updates to the model’s parameters. Thereby, we differentiate between a zero-shot setting (ZS), where the model performs a task based solely on an instruction in the prompt, and few-shot ICL (FS), where a small number of demonstrations is provided in the prompt to guide the model’s behavior. Demonstrations can be retrieved randomly from the dataset or with sophisticated strategies. They can also be static (S) which we report separately.
  • Chain of Thought (CoT): CoT (Wei et al., 2022) refers to a series of subsequent LLM calls, where the (complex) task is broken down into (easier) subtasks or “thoughts” by the LLM itself and answered in a step-by-step manner, rather than jumping to the solution directly. The process also enhances transparency since intermediate steps can be monitored by users.
  • Fine-tuning (FT): In addition to ICL and CoT, which are solely modifications of the prompt, fine-tuning (Gao et al., 2021) modifies the parameters of the LLM, but mostly affects only the last few layers of the neural network, which are usually the most decisive ones for the outcome.
  • Pre-training (PT): Pre-training refers to the initial phase of training the model on a large corpus of text data to learn the structure and patterns of language.
PS-1 — Manual Configuration
This is true if at least one manual configuration step is required (e.g., when a parser is applied to an unseen log type). This includes manual definition of input parameters such as regular expressions, log formats or other essential parameters for different datasets, which is often required for many conventional log parsers such as Drain (He et al., 2017), Brain (Yu et al., 2023), or Spell (Du and Li, 2016), and others featured in the LogPAI log parser repository [^3] https://github.com/logpai/logparser; accessed 9-December-2024 (Zhu et al., 2019). This feature does not include optional parameters that are generic enough to be left unchanged for a new log type.
PS-2 — Retrieval-Augmented Generation
Retrieval-augmented generation, or RAG, is a paradigm where the LLM is provided with information retrieval capabilities. In case of log parsing, most approaches utilize a sampling strategy to include either logs or logs and their templates in the prompt. The LLM should then use the provided context to learn the variability and commonality of logs (Li et al., 2023) or learn parsing directly from log-template pairs. Many approaches create clusters, trees, buckets or other kinds of aggregations from which they sample their demonstrations for ICL and retrieve them based on some kind of similarity measure. We differentiate two cases:
  • Random retrieval (R): The process is a random retrieval of demonstrations from training data.
  • Strategic retrieval (S): There is a refined strategy for retrieving demonstrations from specific data storages (e.g., clusters, lists, etc.).
PS-3 — Caching
Some approaches increase their efficiency by storing parsed logs’ templates in some kind of data structure and only call the LLM if a subsequent log does not fit any of the existing templates. These data structures are mostly tree-like structures or similarity-based clusters.
PS-4 — LLM Usage
The employed LLM is used in at least one of three different ways:
  • Direct parsing (dir): The LLM receives one or more logs directly in the prompt and is asked for the corresponding template.
  • post-processor (post): The LLM is used to post-process the log lines, which is mostly merging similar templates or correcting templates based on new information obtained by subsequent log lines in an online approach.
  • Preprocessor (pre): The LLM functions as a helper in preprocessing, for example, for identifying the timestamp, the log format, or other relevant features.
PS-5 — Template Revision
This feature describes whether the templates are revised in a post-processing step, such as merging similar templates or correcting templates, based on the information obtained from new logs (in an incremental approach). This step can be done with or without the help of LLMs.
R-1 — Evaluation Data
Since the performance of parsers is strongly dependent on the data, we report the different types of datasets or dataset collections used in the papers:
  • LogHub (L): The well-known LogHub (Zhu et al., 2019, 2023) repository is widely used for evaluating log analysis methods. It features $16$ annotated datasets from different domains with $2000$ logs each.
  • Corrected LogHub (CL): This is a version of LogHub with corrected templates. The original templates have been observed to have inconsistencies due to inconsistent labeling styles (Khan et al., 2022).
  • LogHub-2.0 (L2): LogHub-2.0 (Jiang et al., 2024b) is based on Loghub and features $14$ large-scale datasets from various domains with numbers of log in the order of $10^{4}$ to $10^{7}$.
  • Custom (CA/CU): If the parser was evaluated with custom data we report whether it is publicly available (CA) or unavailable (CU).
R-2 — Evaluation Metrics
We report the various evaluation metrics employed to assess the effectiveness of log parsing techniques. Each metric provides a unique perspective on the parsing process, focusing on different aspects of effectiveness:
  • Group Accuracy (GA): A log message is considered correctly parsed if and only if its event template corresponds to the same group of log messages as the ground truth does. GA is then the number of correctly parsed log messages divided by the total number of log messages (Zhu et al., 2019). GA is also known as RandIndex (Rand, 1971).
  • F1-score of Group Accuracy (FGA): $N_{p}$ is the number of templates that are generated by a log parser, and $N_{c}$ the number of templates that are correctly parsed by the log parser. $N_{g}$ is the actual correct number of templates in the ground truth. The precision of the GA (PGA) is then $N_{c}/N_{p}$ and the recall of GA (RGA) is $N_{c}/N_{g}$. Then, FGA is the harmonic mean of PGA and RGA (Jiang et al., 2024b).
  • Parsing Accuracy (PA): A log is considered correctly parsed if and only if all its static text and dynamic variables are correctly identified. PA is then the number of correctly parsed log messages divided by the total number of log messages (Dai et al., 2022). PA is the same as message-level accuracy (MLA) (Liu et al., 2022; Xu et al., 2024).
  • Edit Distance (ED): ED is the minimum number of operations, such as insertions, deletions, or substitutions, needed to convert one string into the other (Nedelkoski et al., 2021).
  • Precision Template Accuracy (PTA): A template is correctly parsed from log messages if and only if it is identical (token-by-token) to the ground-truth template(s) of the log messages. PTA is the ratio of correctly identified templates to the total number of identified templates (Khan et al., 2022).
  • Recall Template Accuracy (RTA): Complementary to PTA, RTA is the ratio of correctly identified templates to the total number of ground-truth templates (Khan et al., 2022).
  • F1-score of Template Accuracy (FTA): Is the harmonic mean of PTA and RTA (Jiang et al., 2024b).
  • Other: If metrics other than the above are used we write “other”.
R-3 — Used Models
A variety of different LLMs are used in the analyzed works. We report only the base models, or the name of the model series, of the used LLMs and only the ones used for the task of log parsing.
R-4 — Code availability
For a potential user, it might be essential that a parser is already implemented, thus we report the availability of code repositories. This feature is true if a link was provided to the implementation and if the link does not lead to an empty repository or a non-existent page, and false otherwise.
R-5 — Preprint
This is true if the reviewed paper is a preprint paper. It is possible that the corresponding code repository to the paper, the results, findings, or interpretations are only preliminary, which is why we report this feature.
4. Literature Review Results
The results of our literature review are aggregated in Table 2. The following subsections discuss the results and findings of the literature review in the context of this table.
Table 2. Categorization of the selected papers by features.
Supervision
Processing mode
Learning / Prompting
Manual config.
RAG
Caching
LLM usage
Template corr.
Datasets
Metrics Models
Code availability
Preprint
General Properties Processing Steps Reproducibility
Approach GP-1 GP-2 GP-3 PS-1 PS-2 PS-3 PS-4 PS-5 R-1 R-2 R-3 R-4 R-5
Astekin et al. (Astekin et al., 2024a) un, sup S ZS, S dir CL GA, ED, other GPT, Claude, Llama
Astekin et al. (Astekin et al., 2024b) un, sup S S dir CL other GPT, Claude, Llama, Mistral
Cui et al. (Cui et al., 2024) (LogEval) un S ZS ? dir L PA, ED GPT, Claude, Llama, Gemini, InternLM, Qwen, AquilaChat, Mistral, Baichuan, DevOps, ChatGLM
Fariha et al. (Fariha et al., 2024) un ? FS? pre L other GPT
Huang et al. (Huang et al., 2024) (LUNAR) un B ZS S dir L2 GA, FGA, PA, FTA GPT ?
Ji et al. (Ji et al., 2024) (SuperLog) sup ? PT, ZS? dir L GA, other Llama, GPT, Qwen, OWL
Jiang et al. (Jiang et al., 2024a) (LILAC) sup S FS S dir L2 GA, FGA, PA, FTA GPT
Karanjai et al. (Karanjai et al., 2024) (LogBabylon) un, sup S FT; FS S dir L2 GA, FGA, PTA, PA, RTA, FTA GPT
Le et al. (Le and Zhang, 2023a) un, sup S ZS, FS R dir CL GA, PA, ED GPT
Liu et al. (Liu et al., 2024) (LogPrompt) un B ZS dir L other GPT, Vicuna
Ma et al. (Ma et al., 2024a) (LLMParser) sup S FT; ZS, FS ? dir CL GA, PA Llama, ChatGLM
Ma et al. (Ma et al., 2024b) (OpenLogParser) un B FS S dir L2 GA, PA Llama, Mistral, Gemma, ChatGLM,
Mehrabi et al. (Mehrabi et al., 2024) sup S FT; ZS, S dir CA PA, ED, FGA Mistral, GPT
Pang et al. (Pang et al., 2024) (ONLA-LLM) un, sup S FT; ZS dir CU other Llama
Pei et al. (Pei et al., 2024) (SelfLog) un, sup B FS S dir L GA, PA, PTA, RTA GPT
Sun et al. (Sun et al., 2023) (Loki) un S ZS dir
Sun et al. (Sun et al., 2024a) (Semirald) un, sup S FS, CoT ? dir L PA GPT
Sun et al. (Sun et al., 2024b) (SemiSMAC-¡T¿) sup S FS, CoT ? dir L GA, PA GPT ?
Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) un B S dir CA other OpenChat, Mistral, WizardLM
Wu et al. (Wu et al., 2024) (AdaParser) un S FS S dir, post L, L2 GA, FGA, PA, FTA GPT, Gemini, Claude, DeepSeek, Qwen
Xiao et al. (Xiao et al., 2024) (LogBatcher) un B ZS S dir L2, CL GA, PA, ED GPT
Xu et al. (Xu et al., 2024) (DivLog) sup S FS S dir L PA, PTA, RTA GPT
Xu et al. (Xu and Gau, 2024) (HELP) un, sup B ZS, S; CoT dir L2 GA, FGA, PA, FTA Claude
Yu et al. (Yu et al., 2024) (LogGenius) un T? ZS pre L, CA PA GPT, Gemma
Zhang et al. (Zhang et al., 2024a) (ECLIPSE) un S ZS ? pre L, CU GA, FGA GPT
Zhang et al. (Zhang et al., 2025) (Lemur) un T FS, CoT post L GA, FGA GPT
Zhi et al. (Zhi et al., 2024) (YALP) un S ZS dir CL GA, PA, ED GPT
Zhou et al. (Zhou et al., 2024) un S ZS, FS/S? ? dir L GA, PA, ED GPT
Zhong et al. (Zhong et al., 2024) (LogParser-LLM) un, sup S FT; ZS, FS S dir, post L, L2 GA, PA, FGA, FTA, other GPT, Llama
4.1. Process Pipeline
After reviewing the paper selection, we were able to develop a process pipeline that summarizes all the approaches in a single flow chart, given in Fig. 2. The dashed arrows and boxes represent optional components while the continuous ones represent essential components. The user and the LLM represent key actors that can act on other components. The arrows pointing away from the user indicate steps that can require manual effort. The arrow from user to the logs represent the manual labeling effort for supervised parsers while the arrow pointing to the preprocessing component represents manual configuration. We provide detailed explanations about the components in the following sections.
Figure 2. LLM-based parsing process pipeline. The dashed arrows and boxes represent optional components while the continuous ones represent essential components. The arrows pointing from the LLM to certain process elements describe where the LLM can be applied.
4.2. Supervision (GP-1)
From the reviewed papers $6$ are fully supervised and $9$ approaches describe a supervised and an unsupervised setting. Supervised approaches require labeled logs for ICL or fine-tuning, but the amount of templates varies from a few guiding seed examples (Zhong et al., 2024) to significant proportions of a dataset’s templates (Jiang et al., 2024a; Xu et al., 2024). The works that fine-tune LLMs (Ma et al., 2024a; Zhong et al., 2024; Pang et al., 2024; Karanjai et al., 2024; Mehrabi et al., 2024) all require logs and their templates for the process which is why we classify these approaches as supervised.
Similar to earlier work (Le and Zhang, 2023b), LILAC (Jiang et al., 2024a) and DivLog (Xu et al., 2024) utilize specialized algorithms to sample from the training data. The objective of the algorithms is to maximize the sample diversity of the labeled logs required for ICL. They call this process candidate sampling. LILAC has a sampling algorithm based on hierarchical clustering while DivLog uses the Determinantal Point Process (DPP) (Kulesza and Taskar, 2012). It is noteworthy that the unsupervised parser LogBatcher (Xiao et al., 2024) also uses DPP, but to maximize the sample diversity within the batches of logs that are prompted to the LLM for direct parsing.
Huang et al. (LUNAR) (Huang et al., 2024) empirically study the average FTA of LILAC (Jiang et al., 2024a) and the BERT-based parser LogPPT (Le and Zhang, 2023b) based on the proportion of labeled logs they receive as input. It was found that FTA exhibited a substantial decline in performance when the proportion of labels was reduced to $5\%$, which, in general, remains a relatively high figure, particularly in light of the typically abundant nature of logs. The finding that the more templates are available, the better the parsers’ performance is little surprising (Jiang et al., 2024a; Huang et al., 2024). However, labeled logs are scarce and require substantial manual effort and expertise (Xiao et al., 2024). All supervised parsers of the selection can technically be operated unsupervised at the cost of performance. This may require modifying the prompt so that it corresponds to a zero-shot setting to not confuse the LLM with announced but absent examples.
4.3. Processing Mode (GP-2)
Log parsers can be categorized based on their processing mode, which determines whether they operate online or offline. The $18$ parsers classified under the “stream” category process logs in real-time as they arrive, making them suitable for online processing. These parsers handle data incrementally, allowing for immediate analysis and response which may be vital for downstream tasks using live monitoring. In contrast, “batch” parsers (Liu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024; Vaarandi and Bahsi, 2024) process logs in fixed-sized chunks, meaning they can function in both online and offline settings, depending on the frequency of batch execution. When processing in batches, it is possible to optimize the process for this batch (Xiao et al., 2024). Lastly, parsers under the category “total” (Zhang et al., 2025; Yu et al., 2024; Huang et al., 2024) analyze an entire log dataset at once, which is inherently an offline approach, as it requires the complete dataset to be available before processing begins. For instance, LUNAR (Huang et al., 2024) and Lemur (Zhang et al., 2025) group the logs based on certain features (e.g. length or most frequent shared oken) into buckets. They then create clusters within that buckets that maximize the sample diversity. The concept of maximizing the sample diversity within certain similarity clusters is also used by LogBatcher (Xiao et al., 2024) but for data batches. Similarly, DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this idea for their sampling algorithm.
4.4. Learning / Prompting (GP-3)
4.4.1. In-Context Learning
ICL (Brown et al., 2020) is a prompt-based learning paradigm. It includes learning from instructions but also from examples (Brown et al., 2020). This makes it a cheap way of learning. All approaches therefore apply ICL in some way. Among others, $16$ works use zero-shot ICL, $14$ employ few-shot ICL, and (Vaarandi and Bahsi, 2024; Mehrabi et al., 2024; Xu and Gau, 2024; Astekin et al., 2024a, b) use static examples. We report static examples separately from the few-shot setting in order to illustrate which approaches use dynamic examples. This in turn implies RAG (PS-2).
4.4.2. Chain of Thought
Lemur (Zhang et al., 2025) uses CoT (Wei et al., 2022) to revise similar templates that may belong to the same event in three dialogue rounds. These steps include revising the structure, the semantics and finally, a decision if the templates should be merged. The other works applying CoT (Sun et al., 2024b, a; Xu and Gau, 2024) use it for direct parsing but do not explain in detail how.
4.4.3. Fine-Tuning and Pre-Training
Compared to ICL, fine-tuning is considered a rather computationally costly task (Xiao et al., 2024) since it modifies the parameters of the models itself. Previous research (Mosbach et al., 2023) found that fine-tuning generally outperforms ICL in effectiveness. From our selection, the works of Ma et al. (LLMParser) (Ma et al., 2024a) and Mehrabi et al.(Mehrabi et al., 2024) support this finding. However, fine-tuning also requires labeled logs. As the default setting for the training, (Mehrabi et al., 2024) used $80\%$ of the available templates and $15$ logs per template, ONLA-LLM (Pang et al., 2024) used $80\%$ of the labeled logs of their custom dataset, LLMParser (Ma et al., 2024a) used $50$ labeled logs per dataset, and (Zhong et al., 2024) used $32$ per dataset. Furthermore, Mehrabi et al. (Mehrabi et al., 2024) and Ma et al. (Ma et al., 2024a) find that their fine-tuned LLMs struggle with new and unseen log types in terms of effectiveness and robustness. This raises the question of whether this is due to overfitting, but this is a task for future research. It remains up to the users to decide whether a costly training phase, requiring considerable quantities of labeled logs, but with improved performance on seen logs, is feasible for their use case.
From our selection only Ji et al. (SuperLog) (Ji et al., 2024) pre-trained a model. Pre-training requires vast amounts of data. To this end, Ji et al. created NLPLog, a comprehensive dataset with a quarter million question-answer pairs on log analysis tasks. The researchers report superior performance not only in parsing but also anomaly detection, failure diagnosis, and log interpretation. They also conducted an evaluation on unseen logs, but they do not report any of the common parsing metrics from Sec. 3.2. This hinders the determination of the practical benefits in settings involving unseen logs. Future research could be conducted on this aspect.
4.5. Manual Configuration (PS-1)
The datasets from LogHub and its descendants (Zhu et al., 2019, 2023; Khan et al., 2022) have been used for the evaluation of log parsing for many years now and it is known from existing literature which regular expressions, log formats, or other parameters have to be used to achieve a certain baseline of performance. Some of the reviewed methods require multiple parameters that are tuned specifically for each dataset and claim to achieve superior performance on the evaluation metrics. These matters may lead to unrealistic expectations for real-life applications, where sometimes no previous knowledge of the versatile format of logs is given (Dai et al., 2023). For instance, Dai et al. have shown that the parsing results of (non-LLM-based) parsers like Drain (He et al., 2017), IPLoM (Makanju et al., 2009), LenMA (Shima, 2016), AEL (Jiang et al., 2008) and Logram (Dai et al., 2022) are strongly influenced by their parameter settings. Of the reviewed studies, $9$ require certain configuration parameters. All of these $9$ require the log format of the logs to extract the log content from the raw logs and (Ma et al., 2024b; Zhang et al., 2025) additionally require regular expressions for preprocessing.
The requirement for manual configuration complicates a parser’s generic application for users and industry. Compared to the above mentioned list of non-LLM-based parsers, the reviewed LLM-based parsers exhibit less (essential) configuration parameters. This indicates that involving the LLM in the parsing process makes parsers more generic, thus reducing the need of human effort.
4.6. RAG (PS-2)
Ten approaches describe a sophisticated retrieval process, one describes random retrieval and for six approaches it is not clear how the approaches apply RAG. RAG is used for demonstration retrieval but also for retrieval of multiple target logs where a set of similar logs is chosen. The latter is only relevant for approaches parsing in batches or the whole dataset at once (GP-2). LogBatcher (Xiao et al., 2024), OpenLogParser (Ma et al., 2024b), LUNAR (Huang et al., 2024) prompt batches of logs that are in the same similarity cluster but maximize the dissimilarity within that cluster to highlight the variabilities and commonalities in logs for the LLM. Demonstrations can either be only logs (Karanjai et al., 2024) or logs and their templates (Xu et al., 2024; Jiang et al., 2024a; Pei et al., 2024; Zhong et al., 2024; Astekin et al., 2024a; Xiao et al., 2024; Ma et al., 2024b; Wu et al., 2024). The templates are either available through being parsed before, yet with no guaranty of correctness, or they origin from candidate sets of labeled logs for supervised settings (Xu et al., 2024; Jiang et al., 2024a; Sun et al., 2024b; Zhong et al., 2024; Pei et al., 2024)(GP-1).
As mentioned in (Ma et al., 2024b) the quality of demonstrations is of high importance for the parsing effectiveness since it can also introduce noise and confuse the LLM. The works (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) support this finding in their ablation studies by comparing the performance of their RAG sampling techniques to random retrieval or static examples, which generally achieves lower scores than strategic retrieval.
Another noteworthy phenomenon observed in generative language models is recency bias which refers to the tendency of the model to give disproportionate weight or attention to information that appears closer to the end of the prompt (Zhao et al., 2021). Consistently, Xu et al. (DivLog) (Xu et al., 2024) observed a significant impact on parsing performance based on the order of the retrieved examples in the prompt, whereby an ascending order of examples by similarity achieved the best results. The ordering of demonstrations in ascending order is adopted by multiple works (Xu et al., 2024; Wu et al., 2024; Xiao et al., 2024; Jiang et al., 2024a).
4.7. Caching (PS-3)
As LLM calls are costly compared to traditional algorithms, many approaches therefore create certain structures like clusters (Sun et al., 2024b; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2025; Jiang et al., 2024a; Karanjai et al., 2024; Huang et al., 2024; Pei et al., 2024), (template) trees (Ma et al., 2024b; Jiang et al., 2024a; Wu et al., 2024; Karanjai et al., 2024; Zhong et al., 2024; Karanjai et al., 2024) or other structures containing logs and their already parsed templates. In case the target log (or logs) matches an existing template (from an already parsed log) or is at least very similar, the parsing algorithm skips the LLM call and returns the corresponding template retrieved from the structure for this log. This self-evolutionary process (as Pei et al. (Pei et al., 2024) call it) is visualized in Fig. 2 by the arrow pointing from the templates to the cache component. This approach can greatly increase efficiency, allowing the usage of parsing with LLMs in settings where fast and cost efficient parsing is crucial, with reported runtimes comparable to runtime efficient parsers, like Drain (He et al., 2017). A parsing cache can also enhance effectiveness by mitigating the instability of results that often occurs with generative language models (Jiang et al., 2024a). Astekin et al. (Astekin et al., 2024b) find that even with temperature 00 some models do not answer deterministically. This emphasizes the usefulness of caches even more. However, it is possible that the first retrieved and stored template for an event type is incorrect, even if it matches some of the incoming logs. Such an event can be counteracted by a template revision step (PS-5). All but one paper (Sun et al., 2023), which described an approach with some form of caching, also described a post-processing step for correcting templates.
Since the total number of templates is drastically less than the number of logs, LLM parsing becomes a scalable approach with caching (Zhong et al., 2024). For example, the LogHub-2.0 datasets (Jiang et al., 2024b) contain more than $50$ million logs but less than $3500$ templates. $14$ approaches apply some form of caching, yet some do not explicitly name it “caching”. The structures used for caching can be efficiently and effectively used for sampling similar logs or relevant parsing demonstrations for ICL with RAG. Caching is not bound to the usage with LLMs and can be built into any other parsing approach.
4.8. LLM Usage (PS-4)
We observed three different ways an LLM can be applied: preprocessing, post-processing and direct parsing, from which we derived the main components of the flow chart in Fig. 2.
Three approaches employ the LLM for preprocessing. ECLIPSE (Zhang et al., 2024a) applies the LLM for the extraction of semantic information from the logs that is then used by an LCS- (longest common subsequence) and entropy-based parsing algorithm. Fariha et al. (Fariha et al., 2024) used an LLM for extracting the regular expressions necessary for subsequent parsing with SPELL (Du and Li, 2016). LogGenius (Yu et al., 2024) uses an LLM to augment log groups with low diversity. The actual parsing task is then left to existing unsupervised parsers, like Drain (He et al., 2017) or SPELL (Du and Li, 2016), for which their parsing effectiveness should be enhanced by the increased diversity of the logs.
Five approaches perform post-processing with the LLM. In all cases, this is done to correct the already parsed logs to yield more accurate templates. How this is performed is explained in Sec 4.9.
Direct parsing is the most straightforward way of parsing with LLMs which is utilized in all but $6$ approaches. There are differences in how the approaches prompt the LLM, yet they all have a similar structure. The prompts these approaches use consist of the following components — note that the components are not strictly separated from each other and can be interlaced:
  • Instruction: Instructs the LLM to parse the log(s) and how. The instruction may also contain rules, such as replacement rules for the timestamp or IP address (Sun et al., 2024b) or special treatment rules for logs concerning exceptions, errors or similar.
  • Context explanations (optional): This part contains explanations about logs or log parsing. For instance, Zhong et al. (Zhong et al., 2024) state to improve the LLM’s capabilities of variable identification by including explanations about variable categories as outlined in (Li et al., 2023). Other approaches (Huang et al., 2024; Sun et al., 2024b; Zhong et al., 2024; Sun et al., 2024a; Astekin et al., 2024a; Xu and Gau, 2024) also provide information about variable types in the prompt. The categorization of the variables can also be beneficial for downstream tasks and the interpretability of the results (Li et al., 2023; Zhong et al., 2024).
  • Context examples (optional): This part either contains multiple logs to illustrate their variabilities and commonalities or logs and their templates as parsing demonstrations. This part can be static, or dynamic using RAG. Providing examples standardizes the LLM’s output format, leading to more stable results (Ma et al., 2024b).
  • Output constraints (optional): This part instructs the LLM how to format the output. Some approaches enforce JSON format, but most approaches determine one or two markers (e.g., often backticks or something like /START and /END) so that in a post-processing step the generated template can be extracted easily with regex after the marker or between markers. This improves output quality, because generative language models sometimes generate unwanted output.
  • Target log(s): Contains the log(s) to be parsed. The approaches (Vaarandi and Bahsi, 2024; Xiao et al., 2024; Huang et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024) provide multiple target logs at once in the query, while the rest provide single logs. Providing multiple logs at once can help the LLM in abstracting variable and static parts from the logs (Pei et al., 2024).
Le et al. (Le and Zhang, 2023a) find that with a simple prompt, where no context is provided nor a detailed instruction is given, GPT-3.5 is hardly able to understand the concept of log parsing. Sun et al. (Sun et al., 2024a) achieve a significantly higher PA ($20\%$ to $60\%$) with a prompt that contains extraction rules for direct parsing (on HDFS and BGL with GPT-3.5 and GPT-4) than without extraction rules.
4.9. Template Revision (PS-5)
In total, $14$ papers revise templates in a post-processing step. All works that correct their templates in a postprocessing step also use a cache (PS-3). The works (Zhang et al., 2025; Zhong et al., 2024; Wu et al., 2024; Liu et al., 2024) employ the LLM to revise templates per prompt: Lemur (Zhang et al., 2025) uses CoT in three dialogue rounds for revising semantically similar templates for potential merge. First, the structure is revised, then the semantics, and in a final round the LLM is asked for a solution based on the first two rounds. LogParser-LLM (Zhong et al., 2024) and LogBabylon (Karanjai et al., 2024) generate a parsing tree from the templates during parsing. In case, a loose match is identified the LLM decides whether the template can be merged or not, leading to an update of the template tree, or the creation of a new tree root node. The works (Liu et al., 2024; Ma et al., 2024b; Wu et al., 2024) re-prompt the logs to the LLM, with the same prompt they were initially parsed, to the LLM if they do not match all the related logs until all logs are matched (Liu et al., 2024) or if a certain number of re-prompts is exceeded (Ma et al., 2024b). AdaParser (Wu et al., 2024) also revises templates if their variables contain certain keywords linked to exceptions, failures or interrupts, which they state should not be abstracted with a wildcard (Wu et al., 2024).
Updating the previously parsed templates can adapt the parser to changes in the computer system and help correct faulty templates. Templates are typically faulty because they have static parts of the log identified as a variable, or variable parts were identified as static. Faulty templates can be identified by high similarity with an existing template (Jiang et al., 2024a; Zhang et al., 2024a; Zhong et al., 2024), if not all logs within a similarity cluster can be matched (Xiao et al., 2024), or if the centroids of log clusters move closer together through newer incoming logs added to clusters (Xu and Gau, 2024). Templates can be merged with templates that partially match with an existing template, by traversing a prefix tree (Jiang et al., 2024a; Karanjai et al., 2024; Zhong et al., 2024), LCS (Zhang et al., 2024a; Zhi et al., 2024), implicitly by reentering the logs into the parsing queue (Xiao et al., 2024), or by monitoring whether multiple paths of a prefix tree join together in a subsequent node. A previously parsed template can also be replaced with a new and more permissive template (Zhi et al., 2024; Vaarandi and Bahsi, 2024). More permissive in this context means that more variables were identified for the newer template.
4.10. Datasets (R-1)
All of the papers in our selection use one of the LogHub versions. The most popular dataset is LogHub (Zhu et al., 2023) used in $14$ of the them. LogHub-2.0 (Jiang et al., 2024b) is used $8$ times, and corrected LogHub (Khan et al., 2022) is used $6$ times. The works (Zhang et al., 2024a; Zhong et al., 2024; Xiao et al., 2024; Wu et al., 2024; Yu et al., 2024) evaluate on two or more datasets, or dataset collections. Zhang et al. (ECLIPSE) (Zhang et al., 2024a) and Pang et al. (Pang et al., 2024) use datasets, other than any of the LogHub ones, that are not publicly available, while (Yu et al., 2024; Mehrabi et al., 2024; Vaarandi and Bahsi, 2024) use custom open-source datasets. The usage of custom datasets along with the datasets from the LogHub versions provides a more comprehensive view on the performance of the parsers. While it is positive to use a variety of datasets for the evaluation, it can also hinder the comparability of the parsing results. For instance, Khan et al. reported performance differences of the conventional parsers on LogHub (Zhu et al., 2023) versus corrected LogHub (Khan et al., 2022). Fu et al. (Fu et al., 2022) and Zhang et al. (ECLIPSE) (Zhang et al., 2024a) showed that many state-of-the-art log parsers perform poorly on their custom datasets while performing well on the LogHub datasets (Zhu et al., 2023).
4.11. Metrics (R-2)
Since the correctness evaluation of templates is not straightforward, there is a strong variety of different evaluation metrics. The set of all used metrics is $\{ED,FGA,FTA,GA,PA,PTA,RTA,\texttt{other}\}$. Except for “other”, these are the traditional metrics that have been widely used (Jiang et al., 2024b; Khan et al., 2022). If we compute the Jaccard Index of the used metrics and all available metrics and average the results we get roughly $0.31$, which means that on average only $31\%$ of the full range of metrics is used (including “other” as a metric). More specifically, $5$ publications use only PA, or PA and GA, $7$ use ED, $9$ publications use at least one metric that is insensitive to class imbalance (FGA, FTA), $8$ publications use metrics other than the traditional ones.
The metrics cover different characteristics of correctness. It is therefore important for evaluating parsers to cover the relevant evaluation aspects with these metrics. For instance, Jiang et al. (Jiang et al., 2024b) proposed using FGA and FTA since they are insensitive to class imbalance, while Khan et al. (Khan et al., 2022) recommend using GA, PA, RTA and PTA to cover all aspects (FTA is the harmonic mean of RTA and PTA). For example, the anomaly detection tool, DeepLog (Du et al., 2017), detects anomalies solely based on the event ID. Each event ID corresponds to a unique template. Consequently, it is indifferent if the parsed templates are correct at the template level as long as they are grouped together correctly, which is indicated by a high score for GA. Since detection algorithms can focus on a variety of data characteristics (Beck et al., 2024), it is important to report metrics that cover these characteristics.
4.12. Model (R-3)
An overview of the LLMs used is given in Table 3. We list only the base models or the name of the model series, since there is an overwhelming variety of different versions and sizes of LLMs. The most widely used model series is GPT. The works (Ma et al., 2024a; Liu et al., 2024; Astekin et al., 2024b; Cui et al., 2024; Zhong et al., 2024; Astekin et al., 2024a; Ma et al., 2024b; Wu et al., 2024; Yu et al., 2024; Mehrabi et al., 2024; Vaarandi and Bahsi, 2024; Ji et al., 2024) evaluate multiple LLMs and compare their performance. Since the papers use different evaluation settings (metrics, datasets) or configurations, it is not straightforward to say which LLM performs best. To this end, we refer to the aforementioned works that compare the results of multiple LLMs and specifically the comprehensive benchmark by Cui et al. (LogEval) (Cui et al., 2024) which evaluates $18$ different LLMs (partly from the same series, but with different sizes).
Table 3. Overview of used LLMs.
Base model / Model series Creator Availability Times used
GPT OpenAI API 23
Llama Meta Open-source 8
Claude Anthropic API 5
Mistral Mistral AI Open-source 5
ChatGLM Tsinghua University Open-source 2
Gemini Google API 2
Gemma Google Open-source 2
Qwen Alibaba Cloud Open-source 2
AquilaChat BAAI Open-source 1
Baichuan Baichuan AI Open-source 1
DeepSeek DeepSeek AI Open-source 1
DevOps-Model CodeFuse Open-source 1
InternLM Shanghai AI Lab Open-source 1
OpenChat OpenChat Team Open-source 1
OWL Camel AI Open-source 1
Vicuna LMSYS Open-source 1
WizardLM Microsoft Open-source 1
4.13. Code (R-4)
From our selection, $16$ papers make their implementation available and provide links to their code repositories. At the time of writing, two of them (Huang et al., 2024; Sun et al., 2024b) lead to an empty or non-existing repository, but consider that both are preprint versions. The rest did not make their code open-source. A study by Olzewski et al. (Olszewski et al., 2023) analyzed the reproducibility and replicability of $750$ machine learning papers and their codebases and datasets from Tier 1 security conferences between the years 2013 and 2022. They found that about $59\%$ of papers did not provide any reproducible artifacts. In our study, $45\%$ of papers do not provide any reproducible artifacts.
4.13.1. Code Quality
A large proportion of the approaches do not provide code and therefore can not be used for the evaluation in Sec. 6. Even with available code, some do not provide the comprehensive functionality to replicate the processes described in their papers and comprehensively correcting the code of others is out of scope for this work. At the time of review (February 14th, 2025) the code of Lemur [^4] https://github.com/zwpride/lemur; accessed 14-February-2025. (Zhang et al., 2025) lacks a script for selecting the previously parsed templates for the subsequent CoT template merging process. The code of LogGenius [^5] https://github.com/huashengyihao/LogGenius; accessed 14-February-2025. (Yu et al., 2024) runs into multiple errors due to missing folders and files, while there is no README file providing instructions. The LogEval repository [^6] https://github.com/LinDuoming/LogEval; accessed 14-Febrary-2025. (Cui et al., 2024) does not provide the scripts for parsing nor instructions. LogPrompt [^7] https://github.com/lunyiliu/LogPrompt; accessed 14-February-2025. (Liu et al., 2024) encounters multiple errors due to simple typos. Furthermore, they describe three different prompt strategies, but it is not clear which they used for their result nor do they explain it in the instruction file. Moreover, they do not provide the functionality to run the parsing process except for the self-prompt setting. SelfLog (Pei et al., 2024) proposes a self-evolutionary approach, in which they “retrieve the most similar historical logs and their templates from the data through an Approximate Nearest Neighbors (ANN) search, serving as the corpus for In-Context Learning (ICL)(Pei et al., 2024). One could think that this means they update the database incrementally after each newly parsed log. However, in the code [^8] https://github.com/CSTCloudOps/SelfLog; accessed 14-February-2025. the script provided for the effectiveness evaluation does not contain the self-evolution functionality. They provide a second file for online parsing that does provide this functionality but it does not work ad hoc due to missing files and missing function parameters. Also, by default they do not use ANN but cosine similarity for retrieval which is somewhere hidden in the code (not an adjustable parameter). The code repository of SuperLog [^9] https://github.com/J-York/SuperLog; accessed 14-Febrary-2025 (Ji et al., 2024) only provides code for LLM training, but not the described framework for parsing. Moreover, we found that many repositories provide sparse instructions, which further impedes reproducibility and application, especially when the execution scripts do not work ad hoc.
These issues appear to be widespread rather than isolated cases. Similar findings were reported in a large-scale study by Trisovic et al. (Trisovic et al., 2022), which examined thousands of replication code repositories containing R files published between 2010 and 2020. Their analysis revealed that a significant majority of these files failed to execute properly, even after code cleaning. Likewise, research by Olzewski et al. (Olszewski et al., 2023) showed that more than half of the artifacts from nearly $300$ reviewed papers could not be run at all. Even among the repositories that did execute, only a fraction produced the expected results, while others either generated different outcomes or lacked crucial components such as arguments or outputs.
4.13.2. Licenses
As we focus on an application perspective, we also report the code licenses for the approaches for which the code is available. They are given in Table 4. Keep in mind that it is possible that code repositories of preprint papers might be preliminary versions or unfinished.
Table 4. Licenses of the approaches. Most approaches do not provide a license. Please note, that especially for preprint papers the license might change or, if absent, be added in the future (status 7-March-2025).
Approach License Preprint?
Cui et al. (Cui et al., 2024) (LogEval) N/A
Ji et al. (Ji et al., 2024) (SuperLog) Apache License Version 2.0, January 2004
Jiang et al. (Jiang et al., 2024a) (LILAC) N/A
Liu et al. (Liu et al., 2024) (LogPrompt) N/A
Ma et al. (Ma et al., 2024a) (LLMParser) N/A
Ma et al. (Ma et al., 2024b) (OpenLogParser) N/A
Mehrabi et al. (Mehrabi et al., 2024) N/A
Pei et al. (Pei et al., 2024) (SelfLog) N/A
Sun et al. (Sun et al., 2024a) (Semirald) N/A
Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) GNU General Public License version 2
Xiao et al. (Xiao et al., 2024) (LogBatcher) MIT License 2024
Xu et al. (Xu et al., 2024) (DivLog) Apache License Version 2.0, January 2004
Yu et al. (Yu et al., 2024) (LogGenius) N/A
Zhang et al. (Zhang et al., 2025) (Lemur) Apache License Version 2.0, January 2004
4.14. Notable Mentions and Promising Techniques
Besides the features we defined in Sec 3.2, we identified other promising techniques and ideas, worth mentioning in the reviewed papers. For instance, some approaches specifically focus on highlighting the commonalities and variabilities of logs. Lemur (Zhang et al., 2025), LUNAR (Huang et al., 2024), OpenLogParser (Ma et al., 2024b), and LogBatcher (Xiao et al., 2024) cluster the logs into multilevel clusters for parsing while the supervised parsers DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this concept for sampling labeled logs. The coarse-grained clusters thereby capture the commonality of the logs, that, in the best case, all belong to the same log group (i.e. the same event), and the fine-grained clusters should highlight the variabilites within that group. This idea is also used by LogShrink (Li et al., 2024) for compressing large log files.
Xiao et al. (LogBatcher) (Xiao et al., 2024) report to use the heuristic rules proposed by Khan et al. (Khan et al., 2022) to correct templates. This includes measures, such as combining subsequent wildcards, like <*><*>, into a single one <*>. This technique is also used by(Xu and Gau, 2024; Jiang et al., 2024a; Xu et al., 2024) and other work outside of our selection (Liu et al., 2022; Le and Zhang, 2023b). By manually inspecting some the templates generated by the parsers used for the evaluation, we found that a significant proportion of templates would have been considered correct by the evaluation functions, if some of these steps would have been applied.
5. Evaluation Setup
In terms of evaluation, the existing literature regarding LLM-based log parsing misses a common ground. While there are some commonalities regarding single features like used datasets, used evaluation metrics, or used models, the combination of these is highly varying between the approaches - see feature R-1, R-2, and R-3 in Table 2. As a consequence, we create a benchmark featuring a subset of $7$ out of the $29$ approaches. To counteract randomness from sampling, LLM output and runtime fluctuations, we run each parser $3$ times take the average.
5.1. Selection for the Benchmark
For the benchmark we only select parsers which made their code publicly available — see feature R-4 in Table 2. We distinguish between approaches that solely focus on methods changing model parameters, like fine-tuning and pre-training, and approaches that apply the LLM as a tool within a framework. Prior studies showed that fine-tuned (Mehrabi et al., 2024) or pre-trained (Ji et al., 2024) models can achieve SotA accuracy scores without complex pre- and post-processing modules for RAG (PS-2), caching (PS-3), and template revision (PS-5). However, it is clear that improved quality of the LLM’s output leads to overall improved output quality of the whole framework. It is therefore meaningful to separate model-centric and model-wrapping approaches. For our evaluation, we focus on the latter and therefore do not include LLMParser (Ma et al., 2024a), SuperLog (Ji et al., 2024) and the work of Mehrabi et al. (Mehrabi et al., 2024) since they do not focus on frameworks but rather on the LLM itself, even tough they provide their code or model, respectively. We also exclude Semirald (Sun et al., 2024a) from the evaluation since its parsing approach consists only of naive queries to an LLM without other processing steps or model training.
In Sec. 4.13.1 we described that some approaches provide code, but do not provide functioning parsers. They are consequently not included in the benchmark (except for LogPrompt).
Given the above-stated matters the remaining approaches featured in the benchmark are LILAC (Jiang et al., 2024a), OpenLogParser (Ma et al., 2024b), LogBatcher (Xiao et al., 2024), SelfLog (Pei et al., 2024), DivLog (Xu et al., 2024), LLM-TD (Vaarandi and Bahsi, 2024), LogPrompt (Liu et al., 2024). To a large extent, this selection represents the state of the art of LLM-based log parsing frameworks, as their approaches cover most of the aspects of this field. Unfortunately, due to the aforementioned selection criteria, this benchmark could not include any work using LLMs as pre- or post-processors.
5.2. Datasets and Baseline
Following previous work (Xiao et al., 2024; Astekin et al., 2024b, a; Zhi et al., 2024; Ma et al., 2024a; Le and Zhang, 2023a, b) we take the corrected version of LogHub from Khan et al. (Khan et al., 2022) as the default dataset for our evaluation. It consists of $16$ datasets with ground-truth templates from different domains such as distributed and supercomputer systems, and server application. Note that we found minor errors in the “Content” column of the corrected version, mostly additional character spaces, that led to errors when evaluating. We exchanged this column (during runtime) with the “Content” column from the original LogHub version (Zhu et al., 2023) (only the templates are different but the content is the same). In Sec. 6.1, we also use the original LogHub version (Zhu et al., 2023) which contains the non-corrected templates. For the runtime evaluation in Sec. 6.2 we use the large-scale datasets from LogHub-2.0 (Jiang et al., 2024b).
Since the LogHub datasets are so widely used for the evaluation of log parsers (see Sec. 4.5) there is a tendency that the parsers work especially well on these datasets but not on others. To simulate a use case closer to a real life example, we extend the $16$ datasets by a custom dataset based on AIT Log Dataset V2 (Landauer et al., 2023b, 2021). The dataset contains $2000$ audit logs (from russellmitchell/gather/intranet_server/logs/audit) with $8$ unique templates. The logs were manually annotated by ourselves, adhering to the style of the templates from the corrected LogHub version (Khan et al., 2022). Audit logs track system activities, user actions, and changes for security, compliance, and troubleshooting. The corresponding files can be found in our code repository.
For the baseline of the evaluation, we select five non-LLM-based parsers from the LogPAI logparser repository [^10] https://github.com/logpai/logparser; accessed 21-March-2025., namely AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023). AEL, SPELL and Drain have been widely used in the log parsing research and their performance has been thoroughly investigated by (Jiang et al., 2024b; Zhang et al., 2023; Khan et al., 2022). Their papers were published in 2008 (AEL), 2016 (SPELL) and 2017 (Drain). ULP and Brain are newer advancements from the years 2022 and 2023. For the baseline parsers the default configuration was used, but consider that the default configuration of all these parsers contains the log format, regular expressions (but empty for some datasets), and two or more other parameters that are specific to each of the LogHub datasets. For the Audit dataset (Landauer et al., 2023b, 2021) the log format “type=<Type> msg=audit(<Time>): <Content>” was used and the regular expression parameter was left empty. The other parameters were simply the parameter configuration that is used most often for the other datasets. The parameters of the configuration were not specifically tuned.
5.3. Evaluation Metrics
Following previous work (Jiang et al., 2024b) we select Group Accuracy (GA), Parsing Accuracy (PA), and F1-score of Template Accuracy (FTA) which are described in Sec. 3.2. There, we also described the Edit Distance (ED) from which we compute the Normalized Edit Distance (NED) (Marzal and Vidal, 1993) following (Xiao et al., 2024).
GA and PA are sensitive to the total number of log messages, which is problematic since most log datasets contain a large number of logs but a much smaller number of unique templates (Jiang et al., 2024b). Therefore, Jiang et al. (Jiang et al., 2024b) proposed to use F-metrics since they are insensitive to class imbalance which is why we use FTA. Furthermore, we use NED since it constitutes a similarity metric between groundtruth and predicted template. The normalization makes NED also a more intuitive indicator for the effectiveness than ED.
5.4. Implementation and Settings
This section describes the implementation details and the settings used.
5.4.1. Settings
The evaluation was performed using three different LLMs of different sizes (small, medium, large):
  1. (1)
    CodeLlama [^11] https://ollama.com/library/codellama; accessed 7-March-2025 (codellama:7b-instruct) is built on top of Llama2 from Meta and has been designed for code-related task. It offers strong performance while requiring significantly fewer parameters than other models. This model is employed to illustrate the performance of a relatively small and efficient LLM ($7$ Billion parameters). Astekin et al. (Astekin et al., 2024a) report that CodeLlama outperforms other LLMs (such as GPT-3.5) in parsing accuracy (PA). In a separate study, Astekin et al. (Astekin et al., 2024b) find that both CodeLlama and GPT-3.5 deliver stable results for the log parsing task. The model was run on a Ubuntu 24.04 LTS server with a 15-core Intel Xeon Gold 6226R CPU and a Tesla V100S (32 GB) GPU via Ollama [^12] https://ollama.com; accessed 14-March-2025.
  2. (2)
    GPT-3.5131313https://platform.openai.com/docs/models/gpt-3.5-turbo; accessed 7-March-2025 (gpt-3.5-turbo-0125) is a widely known closed-source model, reaching outstanding performance in the earlier stages of the LLM hype. It was chosen because it is the most used one from all reviewed papers and it is considered to be medium sized (the size is actually undisclosed but a paper by researchers of Microsoft (Singh et al., 2023) state it is 20 billion parameters). The model is proprietary and was accessed via the OpenAI API.
  3. (3)
    DeepSeek R1141414https://github.com/deepseek-ai/DeepSeek-R1; accessed 25-February-2025 utilizes a novel reinforcement learning approach for model training. It has recently received wide public attention due to its outstanding performance, while being relatively small (but still large) compared to similarly performing models like OpenAI’s o1 model, in combination with its open-source release. This model is used to assess the potential gains in accuracy that can be achieved by utilizing one of the most effective models (671 billion parameters). DeepSeek R1 was called via API from TogetherAI [^15] https://www.together.ai/; accessed 14-March-2025.
The runtime evaluation with the LogHub-2.0 dataset (Jiang et al., 2024b) was run with GPT-3.5 on an Ubuntu 24.04 LTS server with a 32-core AMD EPYC-Milan CPU and 64 GB RAM.
5.4.2. Configuration and Code Changes
For all parsers we set the initial temperature of the LLMs to 00 for a more deterministic output. Note that some parsers increase the temperature in case of a template mismatch or other reasons for some of their LLM calls.
As mentioned in Sec. 4.13.1, there are some issues with some of the parsers’ code which had to be resolved for the benchmark. To ensure the comparability of the output of the parsers, other changes were also necessary to the code of each parser. These changes are kept as small as possible to not disturb the original design or interfere with performance, or at least as little as possible. To ensure model or API compatibility, we updated deprecated versions of the OpenAI (python) package to the newer working ones and we added the functionality to call the OpenAI, Ollama, and TogetherAI API if not given. To have consistent output format, we ensure that the template placeholder symbols are always <*> and that the output is always a CSV file with at least the columns “Content”, and “EventTemplate”. “Content” refers to the log filtered by the log format (i.e. the log message). Since some parsers do not only parse the log content but the entire log, we modify the input to these parsers to only parse the content as well. As found by He et al. (He et al., 2016b), this typically improves the parsing accuracy.
Specific changes were made to DivLog (Xu et al., 2024) which is one of the earliest works on LLM-based log parsing. It does not feature a caching mechanism and therefore, calls the LLM for each log line. Since this is a significant burden in terms of runtime (and cost) and therefore also unrealistic for real-life adaptation, we added a simple cache consisting only of the set of already parsed templates. The cache returns the corresponding template in the event of a match instead of calling the LLM. This slightly affects the accuracy metrics due to improved output consistency. Additionally, we implemented a minor post-processing step where multiple consecutive wildcards were replaced by a single wildcard <*> for efficient matching, to resolve an issue with the generation of infinite consecutive wildcards for some of the logs. This measure slightly increases DivLog’s accuracy.
LogPrompt (Liu et al., 2024) accepts a maximum prompt size parameter which is by default $3000$, corresponding to roughly $25$ logs parsed at once (as they state). Early experiments revealed that this performs extremely poor, hence we set the maximum prompt size to $1000$. This corresponds to roughly $5$ to $10$ logs per prompt, which is in the recommended range of simultaneously parsed logs by Xiao et al. (Xiao et al., 2024). If a template cannot be matched with a log it is again sent to the LLM. We limit the maximum number of re-prompts to $3$ to avoid infinite LLM calls. LogPrompt also required some code corrections, other than typos. For some datasets, LogPrompt runs into an infinite loop due to a faulty template extraction process from the LLM responses. This was fixed by computing the enumeration of the logs instead of extracting it from the prompt again. From the paper it is not clear whether they used ICL, CoT or the self-prompt paradigm for their parsing performance evaluation, yet the parsing functionality is only given for the self-prompt case.
SelfLog (Pei et al., 2024) requires the user to set up an SQL database, yet the main script (run.py) does not contain the self-evolution functionality described in the paper (updating the database with new templates). Therefore, we removed the log-groundtruth template examples part from the prompt and operated the parser as unsupervised and without RAG (no database).
5.4.3. Sampling for Supervised Parsers
From the selected parsers LILAC (Jiang et al., 2024a) and DivLog (Xu et al., 2024) are supervised parsers and require labeled logs to achieve competitive accuracy. In this study, we emphasize minimal human effort and therefore keep the number of available templates $n$ small, namely $n=2$ and $n=4$. Note that the Apache dataset from LogHub (Zhu et al., 2023) contains the smallest number of templates, which is only $6$. For LILAC and DivLog it has been shown that the performance increases with an increasing number of labeled logs available (Jiang et al., 2024a; Xu et al., 2024). This is obvious as the similarity search retrieves the exact template the target log belongs to with a higher likeliness. Multiple works from our selection (Jiang et al., 2024a; Xu et al., 2024) as well as (Le and Zhang, 2023b) state that having a diverse candidate set is crucial to reduce the risk of overfitting to a specific log template, which is why they create specific sampling algorithms to sample from labeled logs. However, having a large proportion of templates at hand is an unrealistic setting in real-life application and requires users to label logs manually (Huang et al., 2024).
To ensure comparability, we exchanged the sampling processes of the supervised parsers DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) with a simple custom sampling approach: from all labeled logs of a dataset we randomly select $n$ (unique) templates. For each of those templates we randomly select a matching log and thus yield a set of log-template pairs which constitutes the candidate set for the supervised parsers. We sample a new set before each run. In the further course of this paper, we indicate the two supervised parsers by the suffix “-$n$”, thus LILAC-$n$, DivLog-$n$, with $n\in[2,4]$.
6. Evaluation Results
This section contains the evaluation results with focus on effectiveness and efficiency.
6.1. Effectiveness
6.1.1. Baseline Performance
The performance of the baseline parsers is visualized in Fig. 3. For some reason, the implementation of ULP (Sedki et al., 2022) outputs the templates in a special format which is why the matching process of the evaluation function does not recognize correctly parsed logs, thus PA and FTA are 00 (GA is not affected).
Figure 3. Performance of the baseline parsers on the corrected LogHub dataset including Audit.
6.1.2. Performance of LLM-based Parsers
The evaluation results of the parsers with the three LLMs on LogHub (Zhu et al., 2023) are given in Fig. 4(a), Fig. 4(b), and Fig. 4(c). Note that we excluded LogPrompt from the evaluation with DeepSeek R1 because the LLM was barely able to extract any templates, resulting in scores of roughly 00 in experimental runs. The LLM did not provide the output in the structured format that was requested. Therefore, the template extraction process, which is based on markers positioned directly before the extracted template, was not able to properly extract the template. This resulted in long inner monologues of DeepSeek R1 and maximum reprompts due to log-template mismatches which would have resulted in an unreasonable financial expense given the cost of the LLM and the poor performance.
(a) Performance with CodeLlama.
(b) Performance with GPT-3.5.
(c) Performance with DeepSeek R1.
Figure 4. Performance of the selected parsers on the corrected LogHub datasets including Audit
In the boxplots of Fig. 4(a), Fig. 4(b), and Fig. 4(c) we can see that the best performance on each metric and for each LLM is achieved by LogBatcher (Xiao et al., 2024), followed by LILAC (Jiang et al., 2024a) for $2$ and $4$ shots. LogBatcher and LILAC clearly outperform the conventional parsers, while the conventional parsers outperform the rest of the LLM-based parsers. In general, the performance of each LLM parser is rather stable across the different models. CodeLlama visibly performs worst, but it is also by far the smallest of the models with 7 billion parameters. Given the size difference of GPT-3.5 and DeepSeek R1, it is surprising that it does not outperform GPT-3.5, given its size. Furthermore, we report the exact numbers for the evaluation with GPT-3.5 in Table 5 to show how the parsers performed on each individual dataset.
Table 5. Performance of the selected parsers with GPT-3.5 on the corrected LogHub datasets including Audit
Android
Apache
BGL
HDFS
HPC
Hadoop
HealthApp
Linux
Mac
OpenSSH
OpenStack
Proxifier
Spark
Thunderbird
Windows
Zookeeper
Audit
Average
Baseline
AEL GA 0.77 1.00 0.96 1.00 0.90 0.87 0.57 0.40 0.76 0.54 0.25 0.97 0.90 0.94 0.69 0.92 0.14 0.74
FTA 0.44 0.50 0.22 0.53 0.46 0.24 0.10 0.27 0.17 0.23 0.08 0.44 0.28 0.20 0.27 0.45 0.00 0.29
PA 0.39 0.69 0.34 0.62 0.68 0.51 0.16 0.17 0.17 0.25 0.02 0.67 0.38 0.04 0.15 0.75 0.00 0.35
NED 0.90 0.92 0.84 0.93 0.97 0.70 0.56 0.74 0.71 0.92 0.60 0.97 0.71 0.71 0.90 0.92 0.38 0.79
Brain GA 0.86 1.00 0.99 1.00 0.94 0.95 1.00 0.35 0.94 1.00 0.49 0.53 1.00 0.97 1.00 0.99 0.21 0.84
FTA 0.37 0.67 0.24 0.80 0.46 0.20 0.40 0.40 0.33 0.23 0.21 0.74 0.44 0.34 0.45 0.57 0.00 0.40
PA 0.24 0.70 0.41 0.96 0.66 0.16 0.25 0.17 0.34 0.28 0.11 0.70 0.38 0.04 0.46 0.78 0.00 0.39
NED 0.84 0.92 0.84 1.00 0.88 0.64 0.88 0.82 0.86 0.95 0.96 0.97 0.96 0.82 0.92 0.94 0.41 0.86
Drain GA 0.83 1.00 0.96 1.00 0.89 0.95 0.78 0.42 0.79 0.79 0.22 0.76 0.92 0.96 1.00 0.97 0.14 0.79
FTA 0.52 0.50 0.22 0.53 0.39 0.30 0.11 0.40 0.20 0.44 0.02 0.41 0.40 0.25 0.41 0.52 0.00 0.33
PA 0.55 0.69 0.34 0.63 0.65 0.51 0.23 0.18 0.22 0.51 0.02 0.68 0.38 0.05 0.46 0.79 0.00 0.41
NED 0.90 0.92 0.84 0.93 0.96 0.79 0.61 0.78 0.76 0.97 0.51 0.96 0.96 0.81 0.86 0.94 0.41 0.82
SPELL GA 0.86 1.00 0.79 1.00 0.65 0.78 0.64 0.15 0.76 0.56 0.26 0.53 0.90 0.84 0.99 0.96 0.14 0.69
FTA 0.25 0.50 0.06 0.43 0.39 0.15 0.10 0.14 0.05 0.23 0.00 0.04 0.19 0.15 0.10 0.37 0.00 0.18
PA 0.15 0.69 0.20 0.30 0.53 0.11 0.15 0.09 0.03 0.19 0.00 0.48 0.32 0.03 0.00 0.75 0.00 0.24
NED 0.81 0.92 0.67 0.94 0.84 0.49 0.52 0.75 0.66 0.86 0.50 0.93 0.67 0.71 0.77 0.94 0.38 0.73
ULP GA 0.74 1.00 0.93 1.00 0.95 0.99 0.90 0.18 0.81 0.43 0.47 0.02 0.92 0.68 0.41 0.99 0.34 0.69
FTA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
PA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
NED 0.63 0.68 0.65 0.58 0.66 0.68 0.74 0.58 0.66 0.75 0.64 0.76 0.67 0.70 0.43 0.76 0.50 0.65
Unsupervised
LLM_TD GA 0.00 1.00 0.50 0.98 0.22 0.47 0.24 0.39 0.00 0.25 0.82 0.27 0.29 0.26 0.57 0.35 0.33 0.41
FTA 0.00 0.50 0.20 0.00 0.19 0.17 0.10 0.14 0.00 0.40 0.57 0.00 0.12 0.04 0.10 0.16 0.00 0.16
PA 0.00 0.69 0.45 0.00 0.20 0.12 0.24 0.28 0.00 0.36 0.25 0.00 0.23 0.10 0.01 0.11 0.00 0.18
NED 0.05 0.95 0.51 0.43 0.26 0.47 0.28 0.42 0.06 0.68 0.46 0.22 0.42 0.29 0.48 0.18 0.09 0.37
LogBatcher GA 0.97 1.00 0.99 1.00 0.95 0.99 1.00 0.84 0.92 1.00 0.98 1.00 1.00 0.90 1.00 0.99 0.27 0.93
FTA 0.78 1.00 0.83 1.00 0.75 0.74 0.95 0.75 0.50 0.82 0.79 1.00 0.76 0.67 0.62 0.81 0.13 0.76
PA 0.83 1.00 0.94 1.00 0.94 0.89 0.99 0.84 0.48 0.97 0.77 1.00 0.97 0.84 0.61 0.99 0.00 0.83
NED 0.92 1.00 0.99 1.00 1.00 0.97 1.00 0.94 0.86 0.99 0.94 1.00 0.98 0.96 0.85 1.00 0.77 0.95
LogPrompt GA 0.13 0.19 0.14 0.00 0.17 0.07 0.06 0.12 0.20 0.00 0.01 0.00 0.01 0.07 0.12 0.45 0.00 0.10
FTA 0.12 0.00 0.03 0.00 0.07 0.08 0.06 0.15 0.09 0.00 0.00 0.00 0.03 0.18 0.05 0.07 0.00 0.06
PA 0.24 0.16 0.38 0.00 0.55 0.15 0.27 0.13 0.23 0.13 0.10 0.00 0.25 0.09 0.33 0.50 0.00 0.21
NED 0.78 0.85 0.83 0.49 0.85 0.70 0.71 0.61 0.77 0.73 0.43 0.54 0.82 0.69 0.78 0.86 0.54 0.70
OpenLogParser GA 0.85 1.00 0.96 0.85 0.93 0.96 0.82 0.42 0.81 0.86 0.57 0.76 0.87 0.94 0.99 0.96 0.14 0.81
FTA 0.32 0.50 0.42 0.00 0.59 0.28 0.46 0.45 0.27 0.65 0.05 0.00 0.41 0.36 0.15 0.41 0.00 0.31
PA 0.24 0.69 0.72 0.00 0.89 0.09 0.66 0.24 0.26 0.69 0.06 0.00 0.36 0.06 0.01 0.48 0.00 0.32
NED 0.17 0.46 0.37 0.27 0.30 0.21 0.26 0.27 0.16 0.39 0.24 0.37 0.21 0.19 0.44 0.27 0.48 0.30
SelfLog GA 0.89 1.00 0.97 1.00 0.89 0.99 0.92 0.29 0.79 0.44 0.44 0.53 0.94 0.70 0.72 0.99 0.14 0.74
FTA 0.39 0.33 0.29 0.00 0.16 0.27 0.48 0.31 0.24 0.17 0.54 0.00 0.38 0.35 0.31 0.24 0.00 0.26
PA 0.46 0.28 0.66 0.00 0.22 0.14 0.55 0.09 0.29 0.15 0.31 0.12 0.76 0.08 0.59 0.24 0.00 0.29
NED 0.76 0.85 0.94 0.86 0.65 0.82 0.89 0.79 0.80 0.85 0.73 0.79 0.93 0.82 0.81 0.80 0.51 0.80
Supervised
DivLog-2 GA 0.56 0.91 0.60 0.49 0.52 0.84 0.68 0.16 0.38 0.47 0.30 0.72 0.46 0.36 0.52 0.93 0.42 0.55
FTA 0.39 0.48 0.32 0.06 0.23 0.31 0.57 0.32 0.19 0.33 0.11 0.39 0.23 0.31 0.19 0.59 0.48 0.32
PA 0.52 0.70 0.72 0.57 0.77 0.32 0.76 0.33 0.26 0.64 0.24 0.76 0.67 0.35 0.31 0.80 0.49 0.54
NED 0.78 0.92 0.93 0.76 0.90 0.80 0.92 0.79 0.52 0.95 0.52 0.95 0.87 0.82 0.70 0.92 0.94 0.82
DivLog-4 GA 0.67 0.53 0.55 0.65 0.36 0.70 0.86 0.20 0.59 0.59 0.27 0.32 0.67 0.40 0.52 0.87 0.36 0.53
FTA 0.42 0.43 0.22 0.25 0.18 0.25 0.56 0.34 0.30 0.55 0.10 0.29 0.48 0.30 0.22 0.53 0.45 0.34
PA 0.50 0.70 0.67 0.71 0.53 0.32 0.76 0.38 0.36 0.65 0.25 0.85 0.77 0.27 0.63 0.79 0.56 0.57
NED 0.78 0.91 0.93 0.87 0.63 0.69 0.91 0.81 0.78 0.95 0.48 0.91 0.93 0.64 0.80 0.93 0.77 0.81
LILAC-4 GA 0.95 1.00 0.98 0.94 0.95 0.97 0.91 0.30 0.81 0.75 0.49 0.95 0.97 0.94 0.79 0.99 1.00 0.86
FTA 0.66 1.00 0.89 0.46 0.75 0.59 0.80 0.70 0.47 0.84 0.83 0.92 0.67 0.56 0.57 0.69 0.88 0.72
PA 0.57 1.00 0.96 0.53 0.92 0.83 0.89 0.42 0.49 0.78 0.43 0.99 0.93 0.73 0.70 0.64 0.78 0.74
NED 0.85 1.00 0.99 0.82 0.99 0.95 0.96 0.92 0.86 0.97 0.90 1.00 0.97 0.93 0.85 0.82 0.99 0.93
LILAC-2 GA 0.96 1.00 0.96 0.94 0.88 0.89 0.91 0.29 0.79 0.63 0.49 0.85 0.92 0.96 0.69 0.99 1.00 0.83
FTA 0.70 1.00 0.80 0.50 0.70 0.57 0.80 0.61 0.48 0.75 0.74 0.77 0.66 0.61 0.57 0.70 0.67 0.68
PA 0.57 1.00 0.94 0.48 0.87 0.70 0.89 0.37 0.48 0.64 0.37 0.92 0.88 0.78 0.52 0.66 0.49 0.68
NED 0.87 1.00 0.98 0.91 0.98 0.90 0.96 0.87 0.87 0.94 0.90 0.99 0.99 0.95 0.77 0.87 0.98 0.93
6.1.3. LLM General Performance
Figure 5 shows the performance for CodeLlama, GPT-3.5 and DeepSeek R1 averaged over all parsers and datasets. One can see that CodeLlama performs worst while GPT-3.5 slightly outperforms DeepSeek R1 except for PA. Consequently, the financial implications of employing DeepSeek R1 do not justify the cost. For instance, the OpenAI API for GPT-3.5 (gpt-3.5-0125) costs $0.5\$$ per $1$ million tokens (input and output) while DeepSeek R1 from the TogetherAI API costs $3\$$ for input and $7\$$ for the output per $1$ million tokens. We found that the reasoning steps of DeepSeek R1 are not always helpful and especially when a structured output is required may confuse the LLM. A possible reason could be that the process of log parsing, not being complex enough, leads to overthinking (Cuadron et al., 2025). However, in (Cuadron et al., 2025) the researchers state that DeepSeek R1 is robust against overthinking. The usefulness of reasoning models for log parsing should therefore be investigated in future research.
Figure 5. Averaged performance of all parsers for CodeLlama, GPT-3.5 and DeepSeek R1.
6.1.4. Performance Difference on LogHub and corrected LogHub
As in the study by Khan et al. (Khan et al., 2022), we report the performance difference between the evaluation on the corrected LogHub datasets (Khan et al., 2022) and the original LogHub datasets (Zhu et al., 2023) (performance of corrected minus performance of original) with GPT-3.5. In Fig. 6 we can see that the highest differences are given for LogBatcher (Xiao et al., 2024), followed by LILAC (Jiang et al., 2024a) for both sample numbers. Naturally, LILAC’s and DivLog’s (Xu et al., 2024) samples were collected and evaluated with the same dataset. That the difference is rather on the positive side indicates that they achieve better performance on the corrected LogHub dataset. Interestingly LILAC was originally evaluated with LogHub-2.0 whose templates correspond to the original LogHub templates and not to the corrected ones (Jiang et al., 2024a). In general, we can see a slight tendency that the parsers output templates that correspond more to the corrected datasets’ templates, which suggests that LLMs “intuitively” prefer the corrected templates’ format.
Figure 6. Performance of the parsers with GPT-3.5 on corrected LogHub minus the performance on the original LogHub.
6.1.5. Performance on the Audit Dataset
An analysis of the performance of all parsers on the Audit dataset reveals that only the supervised parser, LILAC (Jiang et al., 2024a), and to some extent also DivLog (Xu et al., 2024), are able to parse the logs effectively. It should be noted that the configuration of the baseline parsers for Audit was not specifically tuned. It appears that the intended templates’ format of the logs is not straightforward for GPT-3.5 without supervised demonstrations. This finding underscores the efficacy of supervised parsers, demonstrating that a mere two templates are sufficient to attain a reasonable performance level, potentially even a single one demonstrating the preferred template style. Nevertheless, it should be noted that this is merely a solitary example. Future research could establish a more extensive benchmark by incorporating a range of novel and diverse datasets.
Figure 7. Performance of the selected parsers with GPT-3.5 only on the Audit data.
6.2. Efficiency
For the evaluation regarding efficiency we select the well-known datasets HDFS and BGL from LogHub-2.0 (Jiang et al., 2024b). HDFS contains $11.2$ million logs with $46$ unique templates while BGL contains $4.6$ million logs with $320$ unique templates. We excluded DivLog (Xu et al., 2024) and LogPrompt (Liu et al., 2024) from this evaluation since they do not utilize caching. The computation would therefore scale linearly with the number of logs and require multiple million LLM calls.
Figure 8 visualizes the computation time and LLM invocation time of the parsers, Fig. 9 shows the number of LLM calls made. The invocation time for conventional parsers is, of course, zero. Outstanding runtime efficiency is achieved by the conventional parser ULP (Sedki et al., 2022) on both datasets. While LLM-TD is the second fastest and also calls the LLM the least times it is rather the opposite on the BGL dataset. As mentioned in Sec. 4.13.1 and 5.4.2 there were some implementation problems with SelfLog (Pei et al., 2024). They provided a faulty script for the efficiency evaluation which is why we ran the evaluation with the script they designed for the evaluation of the $2000$ log lines version of LogHub (Zhu et al., 2023). However, for both HDFS and BGL the processes got killed and are therefore not featured in the plots.
Figure 8. Computation and invocation time of parsers on HDFS and BGL with GPT-3.5.
Figure 9. Number of LLM calls of parsers on HDFS and BGL with GPT-3.5.
7. Discussion
This section summarizes the findings of the literature review from Sec. 4 and of the evaluation from Sec. 6 by answering the research questions from Sec. 1.
7.1. Answers to Research Questions
RQ1: What are the main advantages and disadvantages of LLM-based log parsing approaches and non-LLM-based approaches?
LLM-based log parsing approaches offer key advantages such as adaptability to diverse log formats, the ability to generalize across different system logs, and robustness to unseen log templates as demonstrated by the evaluation on the Audit dataset (Landauer et al., 2023b, 2021) and other work (Fu et al., 2022). Approaches like the unsupervised parser LogBatcher (Xiao et al., 2024) and the supervised parser LILAC (Jiang et al., 2024a) can handle complex and unstructured logs more effectively than traditional methods. However, they also come with drawbacks, including higher computational costs, potential hallucinations and overthinking for reasoning models (Cuadron et al., 2025), and difficulty in interpreting results due to their black-box nature. In contrast, non-LLM-based approaches, such as Brain (Yu et al., 2023) and ULP (Sedki et al., 2022), tend to be more interpretable, computationally efficient, and easier to deploy but lack the flexibility of LLMs in handling evolving log templates.
RQ2: To what extent do LLM-based log parsing approaches rely on labeled data and manual configuration?
LLM-based log parsing approaches typically require labeled datasets for fine-tuning or evaluation, though some methods leverage self-evolutionary approaches using previously parsed templates to substitute the need for a priori labels. However, with only $2$ and $4$ shots of labeled logs, LILAC (Jiang et al., 2024a) achieves superior performance on the corrected LogHub dataset (Khan et al., 2022) and Audit (Landauer et al., 2023b, 2021), compared to the conventional parsers and other LLM-based parsers, excluding LogBatcher (Xiao et al., 2024). Furthermore, LogBatcher demonstrates impressive performance on both runtime and effectiveness while being unsupervised and only requiring the log format as input parameter. Fine-tuning approaches also report outstanding performance, yet their stronger reliance on labeled logs and computational power and the robust and high performance of LogBatcher and LILAC demonstrate that ICL is a cheaper and performant alternative.
In comparison to the baseline parsers AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023), the LLM-based parsers of our evaluation selection require on average less configuration parameters. This finding is particularly noteworthy in light of the satisfactory performance achieved by LogBatcher (Xiao et al., 2024). This finding suggests that the deployment of LLMs can effectively substitute a substantial proportion of manual effort, thus enhance usability.
RQ3: Which techniques can enhance the efficiency or effectiveness of LLM-based log parsing?
It is evident that ICL and fine-tuning present powerful learning paradigms for LLM-based log parsing. They are not mutually exclusive to each other and can be employed simultaneously to improve parsing accuracy. ICL can be enhanced with RAG or with previously parsed or labeled logs as demonstrations. The ablative studies of reviewed papers (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) unanimously conclude that dynamic demonstrations and sophisticated demonstration selection algorithms improve the accuracy scores. Smart caching solutions have been shown to lead to efficiency improvements, in terms of runtime and monetary expenses, since LLM calls are usually costly. Template revision methods that correct, delete or merge related templates retrospectively can further improve cached templates. This can also help adapt the parser to evolving logs due to updates or other behavioral changes of the computer systems without reconfiguration or retraining. Evaluations show that is also possible to improve the effectiveness of parsers with more powerful LLMs, yet the effective application of reasoning models like DeepSeek R1 should be investigated more closely. It has been shown that variable aware prompts improve the effectiveness in direct parsing with LLMs. Naturally, more explanations in the prompt mean more tokens per parsed log which negatively affects the inference time of LLMs. Short and concise instructions and explanations could thereby alleviate this issue. Furthermore, creating coarse-grained clusters to capture the commonalities of logs and fine-grained clusters to capture the variabilities of logs to support the template extraction process constitutes another promising technique, that can be applied by conventional and LLM-based parsers alike.
RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
Key experiment design choices include the selection of benchmark datasets, evaluation metrics, training methodologies, and hardware specifications. Consistency in log datasets and pre-processing methods is crucial for fair comparisons. For instance, not all parsers require the log format as an input parameter, such as LLM-TD (Vaarandi and Bahsi, 2024) and LUNAR (Huang et al., 2024). LUNAR explicitly includes log headers, such as IP addresses, timestamps, or levels, to grasp the full context of the log. For instance, partitioning logs into groups based on their timestamps is also possible with time intervals (Zhou et al., 2024). Previous research (He et al., 2016b) found that parsing only the extracted log message part, according to the log format, improves effectiveness, but especially LLMs with their generalization capabilities could utilize this extended context.
A set of metrics such as GA, PA, FTA, but also NED (or ED), should be standardized across studies to cover all relevant characteristics of templates, since each metrics maps to specific requirements of specific downstream tasks. The results of our evaluation show high variance between these metrics for different parsers, indicating that the selection covers a variety of aspects. The choice of model hyperparameters also significantly impacts comparability. For example, the researchers of Lemur (Zhang et al., 2025) report near-perfect scores of $0.999$ and $0.996$ for FGA and GA on LogHub (Zhu et al., 2023), yet they require log format, regular expressions, and four other (hyperparameter-tuned) parameters for each dataset. This demonstrates the potential performance levels that can be achieved. However, the number of parameters required is impractical for real-world applications and a comparison to the performance of other parsers is not feasible.
The aforementioned experiment design choices concern all types of parsers. Specific to LLM-based parsers, the employed LLM is also an important factor to consider regarding comparability of evaluation results. As demonstrated, CodeLlama, GPT-3.5, and DeepSeek R1 perform significantly different on the effectiveness scores. Additionally, variations in API versions, underlying model updates, and differences in temperature settings further contribute to comparability issues in performance across studies. This highlights the need for standardized benchmarking practices to ensure fair and reliable comparisons between different LLM-based parsing approaches.
RQ5: To what extent are the presented LLM-based parsing methods accessible and reproducible?
The accessibility and reproducibility of the presented LLM-based parsing methods are significantly hindered by issues related to code availability, documentation, and execution reliability. Many approaches do not provide public code, and even when code is available, it often lacks essential components, such as necessary scripts or instructions, making replication difficult. Several repositories suffer from missing files, incorrect implementations, or incomplete functionality that does not align with the descriptions in their respective papers. Furthermore, discrepancies in the code, such as missing scripts for crucial steps or incorrect default settings, complicate reproducibility. Even when modifications are made to ensure compatibility — such as fixing typos, updating libraries, or implementing minor adjustments to improve consistency — these changes highlight the necessity of external intervention to make the methods functional. The need for such corrections aligns with broader findings from Trisovic et al. (Trisovic et al., 2022) and Olzewski et al. (Olszewski et al., 2023), who observed that a majority of replication code repositories contain errors that prevent immediate execution. Olzewski et al. further found that only a small proportion of the running codebases also produce the claimed results of the related papers (Olszewski et al., 2023). Additionally, some parsers require substantial configuration efforts, further limiting accessibility. Whilst minor improvements, such as adding caching mechanisms to DivLog (Xu et al., 2024) or modifying LogPrompt’s (Liu et al., 2024) prompt size, enhance usability, they also introduce deviations from the original implementations, raising concerns about reproducibility. The discrepancy between our own evaluation results and the reported results raises further concerns about the validity of the publications’ evaluations but naturally also about our own evaluation. Future benchmarks are expected to address these concerns. Finally, while log parsers based on LLM show great potential, their current state significantly hinders accessibility and reliable reproduction without substantial external effort.
7.2. Threats to Validity
We identified two major threats to the validity of the results of our evaluation concerning the adoption of the code and randomness.
7.2.1. Mistakes in Code Adoption
In Sec. 4.13.1 we describe issues with the code of some of the parsers and in Sec. 5.4.2 we describe the changes we made due to the mentioned issues and to ensure transparency. We found that a significant number of parsers do not attain the performance levels claimed in their respective papers. This raises the question whether the observed discrepancies can be attributed to our alterations in their code or the inaccuracy of their evaluation. Alternatively, it is possible that other factors, such as the utilization of different language models, are responsible for the observed variations. However, given that the alterations were predominantly minor, we anticipate that the correctness was not negatively influenced. The code has been made available on GitHub for replication. Should any flaws in the adoption of the parsers be identified, we request to report an issue.
7.2.2. Randomness
The validity of results is threatened by the randomness in LLM output and the randomness introduced by the random sampling for the supervised parsers. To mitigate the randomness in LLMs, we set the LLMs’ initial temperature to 00. The temperature determines the creativity of an LLM. However, even a temperature of 00 does not guarantee a deterministic output (Astekin et al., 2024b). Therefore, we repeated the evaluation three times and computed the average of the measured values.
8. Conclusion
In this work, we conducted a systematic review of LLM-based log parsing approaches, analyzed their strengths and weaknesses in comparison to non-LLM-based methods, and performed an extensive benchmark evaluation to assess their effectiveness, efficiency, and usability. Our findings demonstrate that LLM-based log parsers provide notable advantages, particularly in their adaptability to diverse log formats and their capability to generalize across unseen templates. However, they also come with challenges such as high computational costs, susceptibility to hallucinations, and issues with interpretability.
A key observation is that LLM-based log parsers often reduce the need for manual configuration and labeling, making them more accessible to users with limited domain expertise. Techniques such as in-context learning (ICL), fine-tuning, retrieval-augmented generation (RAG) and template revision can significantly enhance efficiency and accuracy of these methods. Sophisticated dynamic demonstration selection and caching strategies have been shown to significantly impact performance, reducing both LLM inference time and cost. However, despite these improvements, only two out of seven LLM-based parsers, namely LILAC (Jiang et al., 2024a) and LogBatcher (Xiao et al., 2024), are able to outperform the non-LLM-based parsers that constitute our baseline. Furthermore, we identified a lack of reproducibility and comparability, as many implementations lack publicly available code or datasets, comprehensive documentation, or consistency with reported results. Our study highlights the necessity of standardized benchmarking practices for LLM-based log parsing. Differences in experimental setups, dataset preprocessing, and metric selection complicate direct comparisons between methods and impede a meaningful selection for user-specific downstream tasks.
Overall, LLM-based log parsing represents a promising direction for automated log analysis. Techniques like caching, RAG, template revision, especially when used in combination, have clearly proven their value in this field, but further research is required to fully realize the potential of LLM-based log parsing. We hope that our systematic review, benchmark study, and insights contribute to the development of more effective, transparent, and user-friendly log parsing solutions. To facilitate future research and reproducibility, we make our evaluation results and source code publicly available on our GitHub repository.
Acknowledgements.
Funded by the European Union under the European Defence Fund (GA no. 101121403 - NEWSROOM) and under the Horizon Europe Research and Innovation programme (GA no. 101168144 - MIRANDA). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work is co-funded by the Austrian FFG Kiras project ASOC (GA no. FO999905301).

### v2

@@
-SoK: LLM-based Log Parsing
+System Log Parsing with Large Language Models: A Review
 Abstract.
-Log data, generated by software systems, provides crucial insights for tasks like monitoring, root cause analysis, and anomaly detection.
+Log data provides crucial insights for tasks like monitoring, root cause analysis, and anomaly detection.
 Due to the vast volume of logs, automated log parsing is essential to transform semi-structured log messages into structured representations.
-Traditional log parsing techniques often require manual configurations, such as defining log formats or labeling data, which limits scalability and usability.
-Recent advances in large language models (LLMs) have introduced the new research field of LLM-based log parsing, offering potential improvements in automation and adaptability.
-Despite promising results, there is no structured overview of these approaches since this is a relatively new research field with the earliest advances published in late 2023.
-This paper systematically reviews $29$ LLM-based log parsing methods, comparing their capabilities, limitations, and reliance on manual effort.
-We analyze the learning and prompt-engineering paradigms employed, efficiency- and effectiveness-enhancing techniques, and the role of LLMs in the parsing process.
-We aggregate the results of the survey in a large table comprising the characterizing features of LLM-based log parsing approaches and derive the general process of LLM-based log parsing, incorporating all reviewed approaches in a single flow chart.
-Additionally, we benchmark seven open-source LLM-based log parsers on public datasets and critically assess their reproducibility.
-Our findings summarize the advances of this new research field and provide insights for researchers and practitioners seeking efficient and user-friendly log parsing solutions, with all code and results made publicly available for transparency.
-parsing, log data, large language models
-copyright: nonedoi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xpublicationmonth: X
+Recent advances in large language models (LLMs) have introduced the new research field of LLM-based log parsing.
+Despite promising results, there is no structured overview of the approaches in this relatively new research field with the earliest advances published in late 2023.
+This work systematically reviews 29 LLM-based log parsing methods.
+We benchmark seven of them on public datasets and critically assess their comparability and the reproducibility of their reported results.
+Our findings summarize the advances of this new research field, with insights on how to report results, which data sets, metrics and which terminology to use, and which inconsistencies to avoid, with code and results made publicly available for transparency.
+log parsing, log data, large language models, LLMs
+copyright: nonejournalyear: nonacmdoi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xpublicationmonth: Xccs: Computing methodologies Information extractionccs: Software and its engineering Software performance
 1. Introduction
-Log data consists of recorded information generated by logging statements in software applications, systems, devices, or networks during operation.
+Log data consists of recorded information generated by logging facilities in software applications, systems, devices, or networks during operation.
 It provides a chronological record of events and contains valuable insights for various tasks, such as monitoring (Chen et al., 2019; Wang et al., 2021), root cause analysis (He et al., 2021; Notaro et al., 2023), and anomaly detection (Landauer et al., 2023a; He et al., 2016a).
@@
 Log parsing refers to the extraction of structured data from semi-structured logs.
-Parsing the logs into the structured format can simply be done by regular expression matching if the correct log templates are given.
-Most methods therefore formulate log parsing as a log template extraction problem (He et al., 2017; Jiang et al., 2024a).
+Parsing the logs into the structured format can simply be achieved by regular expression matching if the correct log templates are given.
+Most methods therefore formulate log parsing as a log template extraction problem (He et al., 2017; Jiang et al., 2024a; Vaarandi and Bahsi, 2024; Yu et al., 2023; Sedki et al., 2022; Dai et al., 2022).
 Ground-truth templates are defined somewhere in logging statements in the source code and are therefore often not available.
 A wide range of log parsing techniques (Shima, 2016; Jiang et al., 2008; Du and Li, 2016; Sedki et al., 2022; Le and Zhang, 2023b) have therefore been proposed to overcome this issue.
-Many state of the art log parsing methods require manual configuration or other manual actions at some point in the parsing process.
+Many state-of-the-art log parsing methods require manual configuration or other manual actions at some point in the parsing process.
 Approaches may involve labeling of some of the logs from the training set (Jiang et al., 2024a; Xu et al., 2024; Le and Zhang, 2023b), definition of log formats (Li et al., 2023) or regular expressions for preprocessing (Yu et al., 2023; He et al., 2017; Sedki et al., 2022), or other parameters (Yu et al., 2023) that fit the parsing algorithm to the data.
@@
 The reason log parsing methods require manual actions is that humans have both semantic and syntactic understanding (Jiang et al., 2024a), but more importantly, a natural generalization capability and potentially expert knowledge about logs and log parsing.
-Many log parsers have already achieved semantic (Le and Zhang, 2023b; Liu et al., 2022; Jiang et al., 2024a) and syntactic abilities (He et al., 2017; Dai et al., 2022; Yu et al., 2023).
-While syntax-based parsers focus on characteristics like occurrence frequencies, or word or log length, to determine the constant part of logs, semantic-based parsers leverage the semantic differences between static and variable parts of log messages (Jiang et al., 2024a; Xiao et al., 2024).
-Knowledge about logs and how to parse them can be learned by language models, such as BERT (Devlin et al., 2019) or its descendants, by fine-tuning them with logs and their templates (Le and Zhang, 2023b; Tao et al., 2022).
-However, these methods often lack generalization and have to be trained or fine-tuned with labeled logs, preferably with logs of the exact same log type as the target logs (Ma et al., 2024a).
-Generalization arises from pre-training on a large corpus of diverse datasets, allowing models to perform well also on novel tasks and possibly enhanced through fine-tuning (FT) (Gao et al., 2021) or in-context learning (ICL) (Brown et al., 2020).
+Many log parsers have already achieved semantic (Le and Zhang, 2023b; Liu et al., 2022) and syntactic abilities (He et al., 2017; Dai et al., 2022; Yu et al., 2023), but many do not achieve satisfactory performance on data other than LogHub (Zhu et al., 2023) as reported by (Fu et al., 2022; Zhang et al., 2024a; Li et al., 2024a) or require substantial amounts of labeled logs for training their models (Xiao et al., 2024).
+While syntax-based parsers focus on characteristics like occurrence frequencies, or word or log length, to determine the static parts of logs, semantic-based parsers leverage the semantic differences between static and variable parts of log messages (Jiang et al., 2024a; Xiao et al., 2024).
+Knowledge about logs and how to parse them can be learned by language models, such as BERT (Devlin et al., 2019) or its descendants, by fine-tuning them with logs and their templates (Le and Zhang, 2023b; Tao et al., 2022; Liu et al., 2022).
+However, these methods often lack generalization and have to be trained or fine-tuned (Gao et al., 2021) with labeled logs, preferably with logs of the exact same log type as the target logs (Ma et al., 2024a).
+Generalization arises from pretraining on a large corpus of diverse datasets, allowing models to perform well also on novel tasks and possibly enhanced through fine-tuning (Gao et al., 2021) or in-context learning (ICL) (Brown et al., 2020).
 This is where large language models (LLMs) come into play.
@@
 In late 2023, Le and Zhang (Le and Zhang, 2023a) reported notable performance of ChatGPT [^1] https://openai.com/index/chatgpt/; accessed 10-March-2025 in a naive setting, where the LLM was presented individual logs and simply asked for their templates.
-In the meantime, other approaches adopted LLMs for this purpose as well, building sophisticated frameworks that utilize LLMs in various ways to either parse logs directly or enhance log parsing, by adopting learning paradigms for LLMs such as ICL (Brown et al., 2020) or fine-tuning (Gao et al., 2021).
-Given the fact that the number of unique templates is significantly lower than the number of logs and the average rate of newly discovered templates decreases as more logs are processed (Jiang et al., 2024a), some LLM-based parsers (Jiang et al., 2024a) are even able to achieve runtime performance comparable with the fast and well-known parser Drain (He et al., 2017).
-To the best of our knowledge, there is an absence of a structured overview and an objective comparison of papers and methods concerning LLM-based log parsing, as this is a rather novel research field which only recently emerged from the popular research field of generative LLMs.
-So far there is merely a survey paper about LLM-based log analysis techniques by Akhtar et al.
-(Akhtar et al., 2025), which provides a rudimentary overview on using LLMs for log parsing and the techniques employed.
-Consequently, this work aims to undertake a systematic review of the literature to identify the commonalities and variabilities of the existing approaches and to present our gained insights.
-Our main focus is an overview of LLM-based log parsing with emphasis on the users’ perspective — thus, how the LLM is applied and how much manual effort the approaches require.
-The aim is therefore to provide a comprehensive overview of the existing approaches and techniques, enabling anyone interested in (LLM-based) log parsing to quickly and easily identify the most suitable solution for their specific use case.
-The aforementioned goals are formulated as the following research questions:
+In the meantime, other approaches adopted LLMs for this purpose as well, building sophisticated frameworks that utilize LLMs in various ways to enhance log parsing by adopting learning paradigms for LLMs such as ICL (Brown et al., 2020) or fine-tuning (Gao et al., 2021).
+Given the fact that the number of unique templates is significantly lower than the number of logs and the average rate of newly discovered templates decreases as more logs are processed (Jiang et al., 2024a), some LLM-based parsers (Jiang et al., 2024a; Pei et al., 2024) are even able to achieve runtime performance comparable with the fast and well-known parser Drain (He et al., 2017).
+To the best of our knowledge, a structured overview and an objective comparison of papers and methods concerning LLM-based log parsing is currently lacking, as this is a rather novel research field which only recently emerged from the popular research field of generative LLMs.
+So far there is only a preprint survey paper by Akhtar et al.
+(Akhtar et al., 2025), which provides an overview of using LLMs for log analysis, including log parsing.
+Although their discussion on log parsing is concise, it lays a valuable foundation for further innovation and development in this area.
+Consequently, our work aims to undertake a systematic review of the literature, identifying the commonalities and variabilities of the existing approaches, defining common terminologies and reporting schemes, highlighting performance enhancing paradigms, and deriving recommendations for future advancements in this field.
+Our main focus is on investigating LLM-based log parsing methods with emphasis on researchers seeking to apply or develop novel log parsing techniques.
+Thus, additional emphasis is placed on the application of the LLM and the manual effort required for the implementation of the various approaches.
+The aforementioned goals are approached by answering the following research questions:
   • RQ1: What are the main advantages and disadvantages of LLM-based log parsing approaches and non-LLM-based approaches?
@@
   • RQ3: Which techniques can enhance the efficiency or effectiveness of LLM-based log parsing?
-  • RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
+  • RQ4: Which experiment design choices hinder the comparability, and hence which guidelines should be adopted in terms of configuration, datasets used, evaluation metrics, and reporting to make results comparable?
   • RQ5: To what extent are the presented LLM-based parsing methods accessible and reproducible?
-We summarize the main contributions of the present work as follows:
-  • We provide a systematic overview of the existing methods, based on a set of feature definitions that characterize LLM-based log parsing approaches, or the corresponding papers, respectively.
+We summarize the main contributions of this work as follows:
+  • We provide a systematic overview of the existing methods and techniques based on a set of feature definitions that characterize LLM-based log parsing approaches, or the corresponding papers, respectively.
   • We derive the general process of LLM-based log parsing, encompassing each parsing approach of each reviewed paper in a single flow chart.
   • We provide a comprehensive benchmark of seven open-source LLM-based log parsing approaches on open-source datasets.
-  • Based on the literature review and our evaluation results, we answer $5$ research questions and derive further findings that may benefit the future research and design of log parsers.
-  • Finally, we make the source code and all the results of our evaluation publicly available in our GitHub repository [^2] https://github.com/ait-aecid/LLM-log-parsing.git to ensure transparency and reproducibility.
-The remainder is structured as follows: Section 2 describes the background and the core concepts and how we understand them.
-Section 3 outlines the survey method, including the search strategy and the definition of the reviewed features.
+  • Based on the literature review and our evaluation results, we answer $5$ research questions, deriving recommendations for future research and the design of (LLM-based) log parsers.
+  • Finally, we make the source code and all the results of our evaluation publicly available in our GitHub repository (Beck, 2025) to ensure transparency and reproducibility.
+The remainder is structured as follows: Section 2 describes the background and the core concepts.
+Section 3 describes the related work.
+Section 4 outlines the survey method, including the literature search strategy and the definition of the reviewed features.
 Section 5 presents the results of the literature review in a large table and describes the findings of each feature.
@@
 We discuss the findings from literature and evaluation in Sec.
-7 and conclude our paper in Sec.
-8.
-2. Background and Related Work
+8, and conclude our paper in Sec.
+9.
+2. Background
 In this section, we describe the relevant background by means of the core concepts relevant to LLM-based log parsing.
 2.1. Log Parsing
-Log parsing is the foremost step in log analytics and denotes the process of extracting structured data from logs.
-Logs are given as single- or multi-line events and are available in textual form (Landauer et al., 2023a).
-The diverse and often un- or semi-structured nature of logs makes log parsing a difficult task.
-It is the precondition for many log analytics tasks, hence the effectiveness of downstream tasks is therefore inherently dependent on the effectiveness of the parser and also determines the runtime efficiency of the whole process, which is especially important for online (real-time) or incremental approaches, such as live-monitoring for cybersecurity (Jiang et al., 2024a).
-Incremental approaches are designed to handle continuous data streams, updating the parsing model dynamically as new log data arrives.
-This allows for real-time analysis and quicker responses to emerging issues.
+Log parsing is the foremost step in log analytics and denotes the extraction of structured data from logs.
+Logs are given as single- or multi-line events, available in textual form (Landauer et al., 2023a).
+The diverse and un- or semi-structured nature of logs makes log parsing a difficult task.
+It is the precondition for many log analytics tasks, hence the effectiveness of downstream tasks is inherently dependent on the effectiveness of the parser.
+Parsing also determines the runtime efficiency of the whole process, which is especially important for online or incremental approaches, such as live-monitoring for cybersecurity (Jiang et al., 2024a).
+Incremental approaches are designed to handle continuous data streams, updating the parsing model dynamically as new log data arrives, allowing for real-time analysis and quick responses to emerging issues.
 In this work, we specifically focus on template extraction, since the extraction of the variables from the log messages is trivial once the (correct) templates are identified, for example, with regular expression matching.
-The template extraction process is therefore the key component in log parsing.
+The template extraction process is therefore the key component of log parsing.
 2.2. Log Templates
@@
 It typically includes placeholders for essential details like timestamps, log levels (e.g., INFO, ERROR), message content, contextual metadata and more (Li et al., 2023).
-Previous research differs between the unstructured log message part of a log and the structured log headers (He et al., 2016b).
-The structured part is thereby straightforward to extract with regular expressions and it has been shown that parsing only the log message part improves the parsing accuracy (He et al., 2016b).
-Many methods (He et al., 2017; Du and Li, 2016; Xu et al., 2024; Du and Li, 2016), therefore require the log format of a dataset as input parameters.
+Previous research differentiates between the unstructured log message part of a log and the structured log headers (He et al., 2016b).
+The structured part is thereby straightforward to extract with regular expressions.
+It has been shown that parsing only the unstructured log message part improves the parsing accuracy for conventional (non-LLM-based) parsers (He et al., 2016b).
+Many methods (He et al., 2017; Du and Li, 2016; Xu et al., 2024; Du and Li, 2016), therefore require the log format ($\neq$ log template) of a dataset as an input parameter.
 An example of a log template from the Apache dataset (Zhu et al., 2023) is given in Fig.
@@
 The log format parameter for this example would be “[<Time>] [<Level>] <Content>”, whereas “Time” and “Level” refers to log headers and “Content” refers to the log message.
-In the remainder, we understand “log format” as this parameter defining the positions of the log headers and of the content within different log types, or different datasets, respectively.
+In the remainder, we understand the “log format” as this parameter defining the positional indicators of the log headers and of the content within different log types, or different datasets, respectively.
 Figure 1. A simple example of log parsing, from logging statement to log file to parsed log.
@@
 A key enhancement in LLMs is instruction tuning which specializes the pretrained LLMs to follow instructions provided in prompts.
-Instruction tuning enables effective ICL.
+Instruction tuning enables effective in-context learning (ICL).
 Instead of modifying model parameters, ICL allows LLMs to learn from prompts that include instructions, examples, and queries, making them more adaptable to specialized applications (Zhang et al., 2024b).
 The term LLM is only loosely defined, hence, for this work, we define LLMs as generative pretrained transformers (GPTs) such as ChatGPT, Mistral or Qwen.
-Models like BERT (Devlin et al., 2019) (encoder-only) are thereby excluded, as we are only interested in typically larger, generative models that excel in both understanding and generating text, making them suitable for a broader range of tasks.
-Many works included BERT into their definition of LLMs which complicated the manual filtering process after the keyword search.
+Models like BERT (Devlin et al., 2019) (encoder-only) are thereby excluded, as we are only interested in generative models that excel in both understanding and generating text, with increased generalization capabilities compared to BERT, making them suitable for a broader range of tasks.
 2.4. Log Parsing with LLMs
 Large-scale model series like ChatGPT and many more have demonstrated potential in this domain.
-Some approaches fine-tune smaller LLMs (Mehrabi et al., 2024) or BERT (Le and Zhang, 2023b) to this specific task while reporting superior parsing performance to state of the art approaches.
-However, fine-tuning LLMs for log parsing is resource-intensive in terms of computational costs, runtime, as well as labeled examples, making ICL a compelling alternative (Xiao et al., 2024).
+Some approaches fine-tune smaller LLMs (Mehrabi et al., 2024) or BERT (Le and Zhang, 2023b) to this specific task while reporting superior parsing performance to state-of-the-art approaches.
+However, fine-tuning LLMs for log parsing is resource-intensive in terms of computational costs, runtime, and labeling examples, making ICL a compelling alternative (Xiao et al., 2024).
 Research shows that ICL enhances LLMs’ performance in logical reasoning and fact retrieval, suggesting its viability for structured log analysis (Brown et al., 2020).
-Despite these advantages, integrating LLMs into log parsing presents challenges.
-These models are not inherently designed for log parsing, leading to inconsistencies in token classification and errors in log templates.
-Such inaccuracies can negatively impact downstream tasks like anomaly detection (Zhu et al., 2019).
-Additionally, given the vast volume of real-world log data, efficient processing is crucial.
-The computational demands of LLMs contribute to inference overhead and network latency, necessitating optimizations for practical deployment (Jiang et al., 2024a).
+However, LLM inference is also costly.
+The computational demands of LLMs contribute to inference overhead and network latency, but also energy consumption and thus, an increased carbon footprint.
+Using LLMs for automation therefore necessitates optimizations for practical and energy efficient deployment (Jiang et al., 2024a; Xiao et al., 2024).
+Regardless of the efficiency, LLMs are not inherently designed for log parsing, leading to inconsistencies in token classification and errors in log templates.
+Inaccuracies in the parsed templates can negatively impact downstream tasks like anomaly detection (Zhu et al., 2019).
+Despite the aforementioned challenges there is a variety of approaches that address these concerns and provide solutions, making LLM-based log parsing a strong alternative to conventional log parsing.
+3. Related Work
+Existing surveys and benchmarks have significantly contributed to the understanding and evaluation of log parsing.
+These works are particularly relevant to our work, as they provide a comprehensive overview of the current state of the field, highlighting key methodologies, challenges, and advancements: The work of Zhu et al.
+(Zhu et al., 2019) provides a comprehensive benchmark of log parsing techniques, as well as LogHub (Zhu et al., 2023), which became the common dataset collection used in many log analysis evaluations.
+The labels of the LogHub datasets were corrected by Khan et al.
+(Khan et al., 2022) to a unified template style.
+Jiang et al.
+(Jiang et al., 2024b) conducted a large-scale evaluation on state-of-the-art log parsing techniques while also providing the large-scale dataset collection, LogHub-2.0.
+It contains several million annotated logs and has been widely adopted for the evaluation of log analysis techniques (Jiang et al., 2024a; Xiao et al., 2024; Karanjai et al., 2024; Huang et al., 2024; Zhong et al., 2024).
+The large number of logs in LogHub-2.0 facilitated comprehensive efficiency evaluations, which are becoming increasingly important due to the large number of logs generated by modern computer systems (Mi et al., 2013; Zhang et al., 2023).
+Similarly to our work, Zhang et al.
+(Zhang et al., 2023) classify existing parsing approaches by a set of features, such as the parsing mode or the required amount of preprocessing effort.
+The researchers contemplate the challenges of state-of-the-art log parsing techniques, more specifically, the scarcity of public datasets, limited generalizability caused by intrinsic limitations, a lack of automation, and insufficient efficiency regarding the implementation.
+The integration of LLMs into log analysis methods has been explored in the recent survey paper by Akhtar et al.
+(Akhtar et al., 2025).
+It provides a comprehensive survey of LLM-based techniques for log analysis, highlighting their potential to enhance efficiency and accuracy in processing substantial volumes of log data.
+The authors discuss various applications, including anomaly detection, fault monitoring, root cause analysis, threat detection, but also provides an overview log parsing.
+The paper explores the use of fine-tuning, ICL, and retrieval-augmented generation (RAG) (Lewis et al., 2020) for performance improvements in log parsing.
+It also reviews existing research, identifying common challenges such as the high resource consumption of LLMs, the need for frequent updates, and security concerns with closed-source models.
 4. Survey Method
+This section describes the search strategy and the taxonomy of LLM-based log parsing.
 4.1. Search Strategy
-Subsequently, this section describes our search strategy and the inclusion and exclusion criteria derived from the core concepts of Sec.
+This section describes our search strategy and the inclusion and exclusion criteria derived from the core concepts of Sec.
 2 and early insights from LLM-based log parsing approaches.
@@
   • A more recent publication is available that presents the same study (by the same authors), ensuring that this SoK focuses on the latest versions of specific approaches.
-  • It only applies an existing approach without introducing any novelties to LLM-based log parsing.
+  • It only applies a non-LLM-based approach without introducing novelties to LLM-based log parsing.
   • It does not apply log parsing or LLMs by definitions of Sec. 2.
-  • It is not accessible with our institutions’ licenses.
 By applying these exclusion criteria on both abstract and full text, we omitted $147$ of the initial publications found: $1$ publication was omitted because it was not written in English.
+$5$ were outdated versions of newer papers.
+$21$ were omitted because they were books, theses, dissertations, or technical reports.
 $8$ publications were omitted due to being completely unrelated to the topic.
-$110$ are related but excluded, as they do not directly cover LLM-supported log data parsing by the definitions of Sec.
-2 or are about other log analytics tasks, but employ parsing-free methods or existing conventional parsers, such as Drain (He et al., 2017) or SPELL (Du and Li, 2016).
-$21$ were omitted because they were books, theses, dissertations, or technical reports.
-$2$ were excluded for not being available freely with any license from our institutions.
-$5$ were outdated versions of newer papers.
+$112$ are related but excluded, as they do not directly cover LLM-based log parsing by the definitions of Sec.
+2 or focus other log analytics tasks, but employ parsing-free methods or existing conventional parsers, such as Drain (He et al., 2017) or SPELL (Du and Li, 2016).
 The final selection consists of $29$ papers.
 One publication is from the year 2023, one from 2025, while the remaining are all from 2024.
-Since there is a recent hype around the usage of LLMs for log related issues, we include pre-print papers to capture the newest research.
-There are $10$ preprint papers in the final selection ($\sim 34\%$), all from $2024$.
+Since there is a recent hype around the usage of LLMs for log related issues, we include preprint papers to capture the newest research.
+There are $12$ preprint papers in the final selection ($\sim 41\%$).
 4.2. Feature Definition
-This section describes each feature and its classes of Table 2 which we report from the analyzed methods.
-We came up with this specific set of features, since they map to our research questions and pop up significantly often in many of the analyzed works.
-Additionally, we think that these are the most relevant for anyone who wants to apply LLM-based log parsing.
-The features either describe general properties (GP), processing steps (PS), or concern the reproducibility (R) of the approaches.
+This section describes each feature and its classes of Table 2, which we report from the analyzed approaches.
+The features describe either general properties (GP), processing steps (PS), or concern the reproducibility (R) of the approaches.
+Features concerning the general properties of the approaches are typical classifications of machine learning approaches.
+The features we categorize as processing steps are taken from the work of Jiang et al.
+(Jiang et al., 2024a) and Pei at al.
+(Pei et al., 2024).
 For reproducibility, we take inspiration from the work of Olszewski et al.
-(Olszewski et al., 2023) where they analyzed machine learning papers of Tier 1 security conferences for reproducibility.
-The researchers defined a set of questions to determine the reproducibility of research findings.
+(Olszewski et al., 2023) where they analyzed the reproducibility of machine learning papers of Tier 1 security conferences and defined a set of questions to determine the reproducibility of research findings.
+During the review process, we found that these features occur with significant frequencies or with significant variations, validating our selection.
+As illustrated in Table 2, this is evident.
 In Table 2 features are reported with “✓” (true), an empty cell (false) or a descriptive name or abbreviation.
@@
 For instance, one papers provides a link to its code repository but the repository is empty, thus we write “?” for the corresponding feature.
-If a paper does not include the feature, or if the feature is not applicable, we leave it blank.
+If a paper does not include the feature, we leave it blank.
 GP-1 — Supervision
-In log parsing a labeled log is one for which a template, acting as the ground truth for that log, is available. We differentiate approaches based on the requirement for labels:
+In log parsing, a labeled log is one for which a template, acting as the ground truth for that log, is available.
+In machine learning approaches are therefore often classified based on the requirement for labels into supervised (Wu et al., 2024; Liu et al., 2022) or unsupervised (Xiao et al., 2024; He et al., 2017) approaches, thus we report:
   • Supervised parsing (sup) requires at least some log instances of the training set to be labeled.
@@
 GP-2 — Parsing Mode
-Many methods are described as online approaches, but their interpretations vary.
-Some consider online processing to be incremental processing or streaming, while others apply batch-wise processing yet still label their method as online.
+Many methods are described as online approaches (He et al., 2017; Liu et al., 2024), but their interpretations vary.
+Some consider online processing to be incremental processing or streaming, while others apply batch-wise processing, yet still label their method as online.
 Additionally, some works do not specify whether their approach is online or offline.
@@
 As a consequence, we devised the following categories that describe how many logs are processed at once:
-  • Stream (S): The logs are processed one-by-one.
-  • Batch (B): Multiple logs are processed at once, whereby it is possible to apply local (within each batch) optimizations. A batch is significantly smaller than the whole dataset.
+  • Stream (S): The log lines are processed one-by-one.
+  • Batch (B): Multiple log lines are processed at once, whereby it is possible to apply local (within each batch) optimizations. A batch is significantly smaller than the entire dataset.
   • Total (T): The entire dataset is processed at once, whereby it is possible to apply global optimizations to the process.
 GP-3 — Learning / Prompting
-We report four different types of learning or prompt engineering techniques. The type of learning can strongly influence how the prompt is designed which is why we report this in a single feature:
+We identified four different types of learning or prompt engineering techniques. The type of learning can strongly influence how the prompt is designed which is why we report this in a single feature:
   • In-context learning (ZS/FS/S): ICL (Brown et al., 2020) leverages the context provided within a single LLM call to generate responses, adapting to specific needs without any updates to the model’s parameters.
@@
 The process also enhances transparency since intermediate steps can be monitored by users.
-  • Fine-tuning (FT): In addition to ICL and CoT, which are solely modifications of the prompt, fine-tuning (Gao et al., 2021) modifies the parameters of the LLM, but mostly affects only the last few layers of the neural network, which are usually the most decisive ones for the outcome.
-  • Pre-training (PT): Pre-training refers to the initial phase of training the model on a large corpus of text data to learn the structure and patterns of language.
+  • Fine-tuning (FT): In addition to ICL and CoT, which are solely modifications of the prompt, fine-tuning (Gao et al., 2021) modifies the parameters of the LLM, but mostly affects only the last few layers of the neural network, which are the most decisive ones for the outcome.
+  • Pretraining (PT): Pretraining refers to the initial phase of training the model on a large corpus of text data to learn the structure and patterns of language.
 PS-1 — Manual Configuration
@@
 Some approaches increase their efficiency by storing parsed logs’ templates in some kind of data structure and only call the LLM if a subsequent log does not fit any of the existing templates.
-These data structures are mostly tree-like structures or similarity-based clusters.
+These data structures can be tree-like structures, similarity-based clusters, or simply lists.
 PS-4 — LLM Usage
@@
 The original templates have been observed to have inconsistencies due to inconsistent labeling styles (Khan et al., 2022).
-  • LogHub-2.0 (L2): LogHub-2.0 (Jiang et al., 2024b) is based on Loghub and features $14$ large-scale datasets from various domains with numbers of log in the order of $10^{4}$ to $10^{7}$.
+  • LogHub-2.0 (L2): LogHub-2.0 (Jiang et al., 2024b) is based on Loghub and features $14$ large-scale datasets from various domains with numbers of logs in the order of $10^{4}$ to $10^{7}$.
   • Custom (CA/CU): If the parser was evaluated with custom data we report whether it is publicly available (CA) or unavailable (CU).
@@
   • Edit Distance (ED): ED is the minimum number of operations, such as insertions, deletions, or substitutions, needed to convert one string into the other (Nedelkoski et al., 2021).
+In log parsing, it is used to calculate the minimum number of operations required to convert the parsed template into the ground-truth template (Xiao et al., 2024).
   • Precision Template Accuracy (PTA): A template is correctly parsed from log messages if and only if it is identical (token-by-token) to the ground-truth template(s) of the log messages.
@@
   • Other: If metrics other than the above are used we write “other”.
+GA and PA are sensitive to the total number of log messages, which may be problematic since most log datasets contain a large number of logs but a much smaller number of unique templates (Jiang et al., 2024b).
+Therefore, Jiang et al.
+(Jiang et al., 2024b) proposed to use F-metrics since they are insensitive to class imbalance.
 R-3 — Used Models
@@
 Approach GP-1 GP-2 GP-3 PS-1 PS-2 PS-3 PS-4 PS-5 R-1 R-2 R-3 R-4 R-5
-Astekin et al. (Astekin et al., 2024a) un, sup S ZS, S dir CL GA, ED, other GPT, Claude, Llama
+Astekin et al. (Astekin et al., 2024a) un, sup S ZS, S dir CL PA, ED, other GPT, Claude, Llama
 Astekin et al. (Astekin et al., 2024b) un, sup S S dir CL other GPT, Claude, Llama, Mistral
 Cui et al. (Cui et al., 2024) (LogEval) un S ZS ? dir L PA, ED GPT, Claude, Llama, Gemini, InternLM, Qwen, AquilaChat, Mistral, Baichuan, DevOps, ChatGLM
-Fariha et al. (Fariha et al., 2024) un ? FS? pre L other GPT
+Fariha et al. (Fariha et al., 2024) un ? FS? pre L GPT
 Huang et al. (Huang et al., 2024) (LUNAR) un B ZS S dir L2 GA, FGA, PA, FTA GPT ?
-Ji et al. (Ji et al., 2024) (SuperLog) sup ? PT, ZS? dir L GA, other Llama, GPT, Qwen, OWL
+Ji et al. (Ji et al., 2024) (SuperLog) sup ? PT, ZS? dir L GA, FTA Llama, GPT, Qwen, OWL
 Jiang et al. (Jiang et al., 2024a) (LILAC) sup S FS S dir L2 GA, FGA, PA, FTA GPT
@@
 Le et al. (Le and Zhang, 2023a) un, sup S ZS, FS R dir CL GA, PA, ED GPT
-Liu et al. (Liu et al., 2024) (LogPrompt) un B ZS dir L other GPT, Vicuna
+Liu et al. (Liu et al., 2024) (LogPrompt) un B ZS dir L FTA GPT, Vicuna
 Ma et al. (Ma et al., 2024a) (LLMParser) sup S FT; ZS, FS ? dir CL GA, PA Llama, ChatGLM
@@
 Mehrabi et al. (Mehrabi et al., 2024) sup S FT; ZS, S dir CA PA, ED, FGA Mistral, GPT
-Pang et al. (Pang et al., 2024) (ONLA-LLM) un, sup S FT; ZS dir CU other Llama
+Pang et al. (Pang et al., 2024) (ONLA-LLM) un, sup S FT; ZS dir CU PA, PTA, RTA Llama
 Pei et al. (Pei et al., 2024) (SelfLog) un, sup B FS S dir L GA, PA, PTA, RTA GPT
@@
 Sun et al. (Sun et al., 2024b) (SemiSMAC-¡T¿) sup S FS, CoT ? dir L GA, PA GPT ?
-Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) un B S dir CA other OpenChat, Mistral, WizardLM
+Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) un B S dir CA FTA, other OpenChat, Mistral, WizardLM
 Wu et al. (Wu et al., 2024) (AdaParser) un S FS S dir, post L, L2 GA, FGA, PA, FTA GPT, Gemini, Claude, DeepSeek, Qwen
-Xiao et al. (Xiao et al., 2024) (LogBatcher) un B ZS S dir L2, CL GA, PA, ED GPT
+Xiao et al. (Xiao et al., 2024) (LogBatcher) un B ZS S dir L2, CL GA, PA, other GPT
 Xu et al. (Xu et al., 2024) (DivLog) sup S FS S dir L PA, PTA, RTA GPT
@@
 2.
-The dashed arrows and boxes represent optional components while the continuous ones represent essential components.
+The dashed arrows and boxes represent optional actions or components while the continuous ones represent essential actions components.
+Optional components can be skipped.
 The user and the LLM represent key actors that can act on other components.
 The arrows pointing away from the user indicate steps that can require manual effort.
-The arrow from user to the logs represent the manual labeling effort for supervised parsers while the arrow pointing to the preprocessing component represents manual configuration.
+The arrow from user to the logs represent the manual labeling effort for supervised parsers while the arrow pointing to the preprocessing component represents manual configuration (which can also affect other components).
 We provide detailed explanations about the components in the following sections.
@@
 5.2. Supervision (GP-1)
-From the reviewed papers $6$ are fully supervised and $9$ approaches describe a supervised and an unsupervised setting.
-Supervised approaches require labeled logs for ICL or fine-tuning, but the amount of templates varies from a few guiding seed examples (Zhong et al., 2024) to significant proportions of a dataset’s templates (Jiang et al., 2024a; Xu et al., 2024).
+From the reviewed papers, $6$ are fully supervised and $9$ approaches describe a supervised and unsupervised setting.
+Supervised ones require labeled logs for ICL or fine-tuning, but the amount of templates varies from a few guiding seed examples (Zhong et al., 2024) to significant proportions of a dataset’s templates (Jiang et al., 2024a; Xu et al., 2024).
 The works that fine-tune LLMs (Ma et al., 2024a; Zhong et al., 2024; Pang et al., 2024; Karanjai et al., 2024; Mehrabi et al., 2024) all require logs and their templates for the process which is why we classify these approaches as supervised.
@@
 5.3. Processing Mode (GP-2)
-Log parsers can be categorized based on their processing mode, which determines whether they operate online or offline.
 The $18$ parsers classified under the “stream” category process logs in real-time as they arrive, making them suitable for online processing.
-These parsers handle data incrementally, allowing for immediate analysis and response which may be vital for downstream tasks using live monitoring.
-In contrast, “batch” parsers (Liu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024; Vaarandi and Bahsi, 2024) process logs in fixed-sized chunks, meaning they can function in both online and offline settings, depending on the frequency of batch execution.
+On the other hand, “batch” parsers (Liu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024; Vaarandi and Bahsi, 2024) process logs in fixed-sized chunks, meaning they can function in both online and offline settings, depending on the frequency of batch execution.
 When processing in batches, it is possible to optimize the process for this batch (Xiao et al., 2024).
 Lastly, parsers under the category “total” (Zhang et al., 2025; Yu et al., 2024; Huang et al., 2024) analyze an entire log dataset at once, which is inherently an offline approach, as it requires the complete dataset to be available before processing begins.
-For instance, LUNAR (Huang et al., 2024) and Lemur (Zhang et al., 2025) group the logs based on certain features (e.g.
-length or most frequent shared oken) into buckets.
+For instance, LUNAR (Huang et al., 2024) and Lemur (Zhang et al., 2025) group the logs based on certain features (e.g., length or most frequent shared token) into buckets.
 They then create clusters within that buckets that maximize the sample diversity.
@@
 5.4.2. Chain of Thought
-Lemur (Zhang et al., 2025) uses CoT (Wei et al., 2022) to revise similar templates that may belong to the same event in three dialogue rounds.
+Lemur (Zhang et al., 2025) uses CoT (Wei et al., 2022) to revise similar templates, that may belong to the same event, in three dialogue rounds.
 These steps include revising the structure, the semantics and finally, a decision if the templates should be merged.
@@
 From our selection only Ji et al.
-(SuperLog) (Ji et al., 2024) pre-trained a model.
-Pre-training requires vast amounts of data.
+(SuperLog) (Ji et al., 2024) pretrained a model.
+Pretraining requires vast amounts of data.
 To this end, Ji et al.
@@
 They also conducted an evaluation on unseen logs, but they do not report any of the common parsing metrics from Sec.
-3.2.
+4.2.
 This hinders the determination of the practical benefits in settings involving unseen logs.
@@
 have shown that the parsing results of (non-LLM-based) parsers like Drain (He et al., 2017), IPLoM (Makanju et al., 2009), LenMA (Shima, 2016), AEL (Jiang et al., 2008) and Logram (Dai et al., 2022) are strongly influenced by their parameter settings.
-Of the reviewed studies, $9$ require certain configuration parameters.
-All of these $9$ require the log format of the logs to extract the log content from the raw logs and (Ma et al., 2024b; Zhang et al., 2025) additionally require regular expressions for preprocessing.
-The requirement for manual configuration complicates a parser’s generic application for users and industry.
-Compared to the above mentioned list of non-LLM-based parsers, the reviewed LLM-based parsers exhibit less (essential) configuration parameters.
-This indicates that involving the LLM in the parsing process makes parsers more generic, thus reducing the need of human effort.
+Of the reviewed studies, $12$ require certain configuration parameters.
+All of them require the log format of the logs to extract the log content from the raw logs and (Ma et al., 2024b; Zhang et al., 2025) additionally require regular expressions for preprocessing.
 5.6. RAG (PS-2)
@@
 The latter is only relevant for approaches parsing in batches or the whole dataset at once (GP-2).
-LogBatcher (Xiao et al., 2024), OpenLogParser (Ma et al., 2024b), LUNAR (Huang et al., 2024) prompt batches of logs that are in the same similarity cluster but maximize the dissimilarity within that cluster to highlight the variabilities and commonalities in logs for the LLM.
+LogBatcher (Xiao et al., 2024), OpenLogParser (Ma et al., 2024b), and LUNAR (Huang et al., 2024) prompt batches of logs that are in the same similarity cluster but maximize the dissimilarity within that cluster to highlight the variabilities and commonalities in logs for the LLM.
 Demonstrations can either be only logs (Karanjai et al., 2024) or logs and their templates (Xu et al., 2024; Jiang et al., 2024a; Pei et al., 2024; Zhong et al., 2024; Astekin et al., 2024a; Xiao et al., 2024; Ma et al., 2024b; Wu et al., 2024).
@@
 The works (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) support this finding in their ablation studies by comparing the performance of their RAG sampling techniques to random retrieval or static examples, which generally achieves lower scores than strategic retrieval.
-Another noteworthy phenomenon observed in generative language models is recency bias which refers to the tendency of the model to give disproportionate weight or attention to information that appears closer to the end of the prompt (Zhao et al., 2021).
+Another noteworthy phenomenon observed in generative language models is recency bias, which refers to the tendency of the model to give disproportionate weight or attention to information that appears closer to the end of the prompt (Zhao et al., 2021).
 Consistently, Xu et al.
@@
 5.7. Caching (PS-3)
-As LLM calls are costly compared to traditional algorithms, many approaches therefore create certain structures like clusters (Sun et al., 2024b; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2025; Jiang et al., 2024a; Karanjai et al., 2024; Huang et al., 2024; Pei et al., 2024), (template) trees (Ma et al., 2024b; Jiang et al., 2024a; Wu et al., 2024; Karanjai et al., 2024; Zhong et al., 2024; Karanjai et al., 2024) or other structures containing logs and their already parsed templates.
+As LLM calls are costly compared to traditional algorithms, many approaches therefore create certain structures like clusters (Sun et al., 2024b; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2025; Jiang et al., 2024a; Karanjai et al., 2024; Huang et al., 2024; Pei et al., 2024), prefix trees (Ma et al., 2024b; Jiang et al., 2024a; Wu et al., 2024; Karanjai et al., 2024; Zhong et al., 2024; Karanjai et al., 2024) or other structures containing logs and their already parsed templates.
 In case the target log (or logs) matches an existing template (from an already parsed log) or is at least very similar, the parsing algorithm skips the LLM call and returns the corresponding template retrieved from the structure for this log.
-This self-evolutionary process (as Pei et al.
-(Pei et al., 2024) call it) is visualized in Fig.
+This self-evolutionary process (Pei et al., 2024) is visualized in Fig.
 2 by the arrow pointing from the templates to the cache component.
@@
 For example, the LogHub-2.0 datasets (Jiang et al., 2024b) contain more than $50$ million logs but less than $3500$ templates.
-$14$ approaches apply some form of caching, yet some do not explicitly name it “caching”.
+$14$ approaches apply some form of caching, yet some do not explicitly name it “caching” and describe it simply as template matching.
 The structures used for caching can be efficiently and effectively used for sampling similar logs or relevant parsing demonstrations for ICL with RAG.
@@
 The approaches (Vaarandi and Bahsi, 2024; Xiao et al., 2024; Huang et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024) provide multiple target logs at once in the query, while the rest provide single logs.
-Providing multiple logs at once can help the LLM in abstracting variable and static parts from the logs (Pei et al., 2024).
+Providing multiple logs at once can help the LLM in abstracting variable and static parts from the logs (Pei et al., 2024; Xiao et al., 2024).
 Le et al.
@@
 Sun et al.
-(Sun et al., 2024a) achieve a significantly higher PA ($20\%$ to $60\%$) with a prompt that contains extraction rules for direct parsing (on HDFS and BGL with GPT-3.5 and GPT-4) than without extraction rules.
+(Sun et al., 2024a) achieve a significantly higher PA ($20\%$ to $60\%$) with a prompt that contains extraction rules for direct parsing (on the HDFS and BGL datasets from LogHub (Zhu et al., 2023) with GPT-3.5 and GPT-4) than without extraction rules.
 5.9. Template Revision (PS-5)
@@
 LogParser-LLM (Zhong et al., 2024) and LogBabylon (Karanjai et al., 2024) generate a parsing tree from the templates during parsing.
-In case, a loose match is identified the LLM decides whether the template can be merged or not, leading to an update of the template tree, or the creation of a new tree root node.
+In case, a loose match is identified the LLM decides if the template should be merged, leading to an update of the template tree, or the creation of a new tree root node.
 The works (Liu et al., 2024; Ma et al., 2024b; Wu et al., 2024) re-prompt the logs to the LLM, with the same prompt they were initially parsed, to the LLM if they do not match all the related logs until all logs are matched (Liu et al., 2024) or if a certain number of re-prompts is exceeded (Ma et al., 2024b).
@@
 Faulty templates can be identified by high similarity with an existing template (Jiang et al., 2024a; Zhang et al., 2024a; Zhong et al., 2024), if not all logs within a similarity cluster can be matched (Xiao et al., 2024), or if the centroids of log clusters move closer together through newer incoming logs added to clusters (Xu and Gau, 2024).
-Templates can be merged with templates that partially match with an existing template, by traversing a prefix tree (Jiang et al., 2024a; Karanjai et al., 2024; Zhong et al., 2024), LCS (Zhang et al., 2024a; Zhi et al., 2024), implicitly by reentering the logs into the parsing queue (Xiao et al., 2024), or by monitoring whether multiple paths of a prefix tree join together in a subsequent node.
+Templates can be merged with templates that partially match with an existing template, by traversing a prefix tree (Jiang et al., 2024a; Karanjai et al., 2024; Zhong et al., 2024), LCS (Zhang et al., 2024a; Zhi et al., 2024), implicitly by reentering the logs into the parsing queue (Xiao et al., 2024), or by monitoring whether multiple paths (templates) of a prefix tree join together in a subsequent node (Pei et al., 2024).
 A previously parsed template can also be replaced with a new and more permissive template (Zhi et al., 2024; Vaarandi and Bahsi, 2024).
@@
 (ECLIPSE) (Zhang et al., 2024a) showed that many state-of-the-art log parsers perform poorly on their custom datasets while performing well on the LogHub datasets (Zhu et al., 2023).
+Using custom datasets alongside LogHub datasets offers a broader view of parser performance but can affect result comparability.
+The studies (Khan et al., 2022; Fu et al., 2022; Zhang et al., 2024a) show that parsers often perform well on LogHub datasets (Zhu et al., 2023) or its descendants (Jiang et al., 2008; Khan et al., 2022) but poorly on custom datasets.
+Additionally, it is not clear whether the open-source log datasets used are known to the LLMs from their extensive pretraining phases.
+However, the ablative studies of (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024) report significant performance improvements with ICL compared to direct zero-shot parsing, arguing that this implies a low likelihood of memorization of single log templates for GPT-3 and GPT-3.5-Turbo.
 5.11. Metrics (R-2)
@@
 Except for “other”, these are the traditional metrics that have been widely used (Jiang et al., 2024b; Khan et al., 2022).
-If we compute the Jaccard Index of the used metrics and all available metrics and average the results we get roughly $0.31$, which means that on average only $31\%$ of the full range of metrics is used (including “other” as a metric).
-More specifically, $5$ publications use only PA, or PA and GA, $7$ use ED, $9$ publications use at least one metric that is insensitive to class imbalance (FGA, FTA), $8$ publications use metrics other than the traditional ones.
+If we compute the Jaccard Index of the used metrics and all available metrics and average the results we get roughly $0.33$, which means that on average only $33\%$ of the full range of metrics is used (including “other” as a metric).
+More specifically, $5$ publications use only PA, or PA and GA, $6$ use ED, $12$ publications use at least one metric that is insensitive to class imbalance (FGA, FTA).
+Also, $5$ publications use metrics other than traditional ones.
+For example, the creators of LogBatcher (Xiao et al., 2024) report the normalized Edit Distance (NED) (Marzal and Vidal, 1993) which computes the mean ED of all template pairs compared in the data set.
+In (Astekin et al., 2024b) Astekin et al.
+report the number of unique templates as a means of LLM determinism, while in (Astekin et al., 2024a) they report variations of metics based on ED and LCS.
+The creators of LogParser-LLM (Zhong et al., 2024) propose metrics based on Granularity Distance, the minimum number of operations necessary to transform one parsing result into the other.
+An operation is thereby defined as converting a static part of the parsed template to a variable or vice versa.
 The metrics cover different characteristics of correctness.
@@
 Consequently, it is indifferent if the parsed templates are correct at the template level as long as they are grouped together correctly, which is indicated by a high score for GA.
-Since detection algorithms can focus on a variety of data characteristics (Beck et al., 2024), it is important to report metrics that cover these characteristics.
+Since downstream tasks, like anomaly detection, can focus on a variety of data characteristics (Beck et al., 2024), it is important to report metrics that cover these characteristics.
+Previous work (Petrescu et al., 2023; Liu et al., 2022; Nedelkoski et al., 2021) found that the state-of-the-art log parsers are often evaluated with insufficient selections of evaluation metrics.
+While their work concerns the conventional non-LLM based approaches, we find that many LLM-based approaches carry on with this manner.
+More specifically, we consider using only PA or the combination of the grouping metrics GA and FGA as insufficient.
+The highly varying scores between the metrics of different parsers in (Jiang et al., 2024b; Khan et al., 2022) but also in our own evaluation in Sec.
+7 show how important a versatile evaluation metric selection is.
+Focusing only on a single aspect misleads estimations of a parser’s performance.
 5.12. Model (R-3)
@@
 (LogEval) (Cui et al., 2024) which evaluates $18$ different LLMs (partly from the same series, but with different sizes).
-Table 3. Overview of used LLMs.
+Table 3. Overview of LLMs used.
 Base model / Model series Creator Availability Times used
@@
 (Olszewski et al., 2023) analyzed the reproducibility and replicability of $750$ machine learning papers and their codebases and datasets from Tier 1 security conferences between the years 2013 and 2022.
-They found that about $59\%$ of papers did not provide any reproducible artifacts.
-In our study, $45\%$ of papers do not provide any reproducible artifacts.
+They found that about $59\%$ of the papers did not provide any reproducible artifacts.
+In our study, $45\%$ of the reviewed papers do not provide any reproducible artifacts (including preprints).
 5.13.1. Code Quality
 A large proportion of the approaches do not provide code and therefore can not be used for the evaluation in Sec.
-6.
+7.
 Even with available code, some do not provide the comprehensive functionality to replicate the processes described in their papers and comprehensively correcting the code of others is out of scope for this work.
@@
 LogPrompt [^6] https://github.com/lunyiliu/LogPrompt; accessed 14-February-2025.
-(Liu et al., 2024) encounters multiple errors due to simple typos.
+(Liu et al., 2024) encounters multiple errors due to simple typos (e.g.
+a variable that seems to be undefined but it is actually spelled differently than in its definition).
 Furthermore, they describe three different prompt strategies, but it is not clear which they used for their result nor do they explain it in the instruction file.
@@
 Table 4.
-Licenses of the approaches.
+Licenses of the approaches with open-source implementations.
 Most approaches do not provide a license.
@@
 Zhang et al. (Zhang et al., 2025) (Lemur) Apache License Version 2.0, January 2004
-4.14. Notable Mentions and Promising Techniques
+5.14. Notable Mentions
 Besides the features we defined in Sec 4.2, we identified other promising techniques and ideas, worth mentioning in the reviewed papers.
@@
 the same event), and the fine-grained clusters should highlight the variabilites within that group.
-This idea is also used by LogShrink (Li et al., 2024) for compressing large log files.
+This idea is also used by LogShrink (Li et al., 2024b) for compressing large log files.
 Xiao et al.
@@
 This includes measures, such as combining subsequent wildcards, like <*><*>, into a single one <*>.
-This technique is also used by(Xu and Gau, 2024; Jiang et al., 2024a; Xu et al., 2024) and other work outside of our selection (Liu et al., 2022; Le and Zhang, 2023b).
+This technique, or similar, is also used by (Xu and Gau, 2024; Jiang et al., 2024a; Xu et al., 2024) and other work outside of our selection (Liu et al., 2022; Le and Zhang, 2023b).
 By manually inspecting some the templates generated by the parsers used for the evaluation, we found that a significant proportion of templates would have been considered correct by the evaluation functions, if some of these steps would have been applied.
@@
 In terms of evaluation, the existing literature regarding LLM-based log parsing misses a common ground.
-While there are some commonalities regarding single features like used datasets, used evaluation metrics, or used models, the combination of these is highly varying between the approaches - see feature R-1, R-2, and R-3 in Table 2.
-As a consequence, we create a benchmark featuring a subset of $7$ out of the $29$ approaches.
-To counteract randomness from sampling, LLM output and runtime fluctuations, we run each parser $3$ times take the average.
+While there are some commonalities regarding single features, such as used datasets, used evaluation metrics, or used models, the combination of these is highly varying between the approaches - see feature R-1, R-2, and R-3 in Table 2.
+As a consequence, we create a benchmark comparing a subset of $7$ out of the $29$ approaches.
+To counteract randomness from sampling, LLM output and runtime fluctuations, we run each parser $3$ times and take the average.
 6.1. Selection for the Benchmark
 For the benchmark we only select parsers which made their code publicly available — see feature R-4 in Table 2.
-We distinguish between approaches that solely focus on methods changing model parameters, like fine-tuning and pre-training, and approaches that apply the LLM as a tool within a framework.
-Prior studies showed that fine-tuned (Mehrabi et al., 2024) or pre-trained (Ji et al., 2024) models can achieve SotA accuracy scores without complex pre- and post-processing modules for RAG (PS-2), caching (PS-3), and template revision (PS-5).
+We distinguish between approaches that solely focus on methods changing model parameters, like fine-tuning and pretraining, and approaches that apply the LLM as a tool within a framework.
+Prior studies showed that fine-tuned (Mehrabi et al., 2024) or pretrained (Ji et al., 2024) models can achieve state-of-the-art accuracy scores without pre- and post-processing modules for RAG (PS-2), caching (PS-3), and template revision (PS-5).
 However, it is clear that improved quality of the LLM’s output leads to overall improved output quality of the whole framework.
-It is therefore meaningful to separate model-centric and model-wrapping approaches.
-For our evaluation, we focus on the latter and therefore do not include LLMParser (Ma et al., 2024a), SuperLog (Ji et al., 2024) and the work of Mehrabi et al.
-(Mehrabi et al., 2024) since they do not focus on frameworks but rather on the LLM itself, even tough they provide their code or model, respectively.
-We also exclude Semirald (Sun et al., 2024a) from the evaluation since its parsing approach consists only of naive queries to an LLM without other processing steps or model training.
-In Sec. 4.13.1 we described that some approaches provide code, but do not provide functioning parsers. They are consequently not included in the benchmark (except for LogPrompt).
+It is therefore meaningful to separate model-centric and model-wrapping approaches — especially, in the rapidly evolving landscape of LLMs, where more performant models are introduced on a daily basis.
+For our evaluation, we limit our scope and focus on these frameworks and therefore do not include the fine-tuning or pretraining approaches LLMParser (Ma et al., 2024a), SuperLog (Ji et al., 2024), and the work of Mehrabi et al.
+(Mehrabi et al., 2024), even tough they provide their code or model.
+We also exclude Semirald (Sun et al., 2024a) since it does not provide a log parsing framework and focuses rather on anomaly detection.
+In Sec.
+5.13.1, we described that some approaches provide code, but do not provide functioning parsers.
+They are consequently not included in the benchmark (except for LogPrompt since its typos could be resolved by exchanging single characters).
 Given the above-stated matters the remaining approaches featured in the benchmark are LILAC (Jiang et al., 2024a), OpenLogParser (Ma et al., 2024b), LogBatcher (Xiao et al., 2024), SelfLog (Pei et al., 2024), DivLog (Xu et al., 2024), LLM-TD (Vaarandi and Bahsi, 2024), LogPrompt (Liu et al., 2024).
-To a large extent, this selection represents the state of the art of LLM-based log parsing frameworks, as their approaches cover most of the aspects of this field.
+To a large extent, this selection represents the state-of-the-art of LLM-based log parsing frameworks, as their approaches cover most of the aspects of this field.
 Unfortunately, due to the aforementioned selection criteria, this benchmark could not include any work using LLMs as pre- or post-processors.
@@
 In Sec.
-6.1, we also use the original LogHub version (Zhu et al., 2023) which contains the non-corrected templates.
+7.1, we also use the original LogHub version (Zhu et al., 2023) which contains the non-corrected templates.
 For the runtime evaluation in Sec.
-6.2 we use the large-scale datasets from LogHub-2.0 (Jiang et al., 2024b).
+7.2 we follow previous work (Jiang et al., 2024a) and only report the runtime of the large-scale datasets from LogHub-2.0 (Jiang et al., 2024b).
+This choice is motivated by the need for a more extensive data as the $2000$ samples from the original LogHub are insufficient for a comprehensive runtime analysis.
+The results for the runtime on the LogHub datasets are available at (Beck, 2025).
 Since the LogHub datasets are so widely used for the evaluation of log parsers (see Sec.
-4.5) there is a tendency that the parsers work especially well on these datasets but not on others.
+5.5) there is a tendency that the parsers work especially well on these datasets but not on others.
 To simulate a use case closer to a real life example, we extend the $16$ datasets by a custom dataset based on AIT Log Dataset V2 (Landauer et al., 2023b, 2021).
 The dataset contains $2000$ audit logs (from russellmitchell/gather/intranet_server/logs/audit) with $8$ unique templates.
+This dataset will be referred to as Audit in the following.
 The logs were manually annotated by ourselves, adhering to the style of the templates from the corrected LogHub version (Khan et al., 2022).
 Audit logs track system activities, user actions, and changes for security, compliance, and troubleshooting.
-The corresponding files can be found in our code repository.
+The corresponding files can be found in our code repository (Beck, 2025).
 For the baseline of the evaluation, we select five non-LLM-based parsers from the LogPAI logparser repository [^9] https://github.com/logpai/logparser; accessed 21-March-2025., namely AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023).
@@
 ULP and Brain are newer advancements from the years 2022 and 2023.
-For the baseline parsers the default configuration was used, but consider that the default configuration of all these parsers contains the log format, regular expressions (but empty for some datasets), and two or more other parameters that are specific to each of the LogHub datasets.
+The baseline parsers use the standard configuration of the literature (Xiao et al., 2024; Jiang et al., 2024a; Zhu et al., 2023; Jiang et al., 2024b).
+The default configuration of all these parsers contains the log format, regular expressions (but empty for some datasets), and up to two other parameters that are specific to each of the LogHub datasets.
 For the Audit dataset (Landauer et al., 2023b, 2021) the log format “type=<Type> msg=audit(<Time>): <Content>” was used and the regular expression parameter was left empty.
-The other parameters were simply the parameter configuration that is used most often for the other datasets.
-The parameters of the configuration were not specifically tuned.
+The other parameters for the Audit dataset were chosen through grid search by the best score for GA.
+We used the full dataset for this hyperparameter tuning to show the best possible performance of the baseline.
 6.3. Evaluation Metrics
-Following previous work (Jiang et al., 2024b) we select Group Accuracy (GA), Parsing Accuracy (PA), and F1-score of Template Accuracy (FTA) which are described in Sec.
-3.2.
-There, we also described the Edit Distance (ED) from which we compute the Normalized Edit Distance (NED) (Marzal and Vidal, 1993) following (Xiao et al., 2024).
-GA and PA are sensitive to the total number of log messages, which is problematic since most log datasets contain a large number of logs but a much smaller number of unique templates (Jiang et al., 2024b).
-Therefore, Jiang et al.
-(Jiang et al., 2024b) proposed to use F-metrics since they are insensitive to class imbalance which is why we use FTA.
-Furthermore, we use NED since it constitutes a similarity metric between groundtruth and predicted template.
-The normalization makes NED also a more intuitive indicator for the effectiveness than ED.
+Following previous work (Jiang et al., 2024b) we report Group Accuracy (GA), Parsing Accuracy (PA), and F1-score of Template Accuracy (FTA) which are described in Sec.
+4.2.
+GA and PA are the most used ones in the literature (cf.
+Sec.
+5.11) and FTA is recommended by Joang et al.
+(Jiang et al., 2024b).
+Section 4.2 also described the Edit Distance (ED), which is the recommended parsing metric of Petrescu et al.
+(Petrescu et al., 2023), from which we compute the Normalized Edit Distance (NED) (Marzal and Vidal, 1993).
+From our selection, NED is only used by Xiao et al.
+(LogBatcher) (Xiao et al., 2024).
+Nevertheless, we report NED because normalization makes it a more intuitive indicator of effectiveness than ED, as it fits into the $[0,1]$ range of the other conventional metrics.
 6.4. Implementation and Settings
@@
 6.4.1. Settings
-The evaluation was performed using three different LLMs of different sizes (small, medium, large):
+The evaluation was performed using three LLMs of different sizes (small, medium, large):
   1. (1)
@@
   2. (2)
-    GPT-3.5131313https://platform.openai.com/docs/models/gpt-3.5-turbo; accessed 7-March-2025 (gpt-3.5-turbo-0125) is a widely known closed-source model, reaching outstanding performance in the earlier stages of the LLM hype.
+    GPT-3.5121212https://platform.openai.com/docs/models/gpt-3.5-turbo; accessed 7-March-2025 (gpt-3.5-turbo-0125) is a widely known closed-source model, reaching outstanding performance in the earlier stages of the LLM hype.
 It was chosen because it is the most used one from all reviewed papers and it is considered to be medium sized (the size is actually undisclosed but a paper by researchers of Microsoft (Singh et al., 2023) state it is 20 billion parameters).
@@
   3. (3)
-    DeepSeek R1141414https://github.com/deepseek-ai/DeepSeek-R1; accessed 25-February-2025 utilizes a novel reinforcement learning approach for model training.
+    DeepSeek R1131313https://github.com/deepseek-ai/DeepSeek-R1; accessed 25-February-2025 utilizes a novel reinforcement learning approach for model training.
 It has recently received wide public attention due to its outstanding performance, while being relatively small (but still large) compared to similarly performing models like OpenAI’s o1 model, in combination with its open-source release.
@@
 DeepSeek R1 was called via API from TogetherAI [^14] https://www.together.ai/; accessed 14-March-2025.
-The runtime evaluation with the LogHub-2.0 dataset (Jiang et al., 2024b) was run with GPT-3.5 on an Ubuntu 24.04 LTS server with a 32-core AMD EPYC-Milan CPU and 64 GB RAM.
+The runtime evaluation with the LogHub-2.0 dataset (Jiang et al., 2024b) was run on an Ubuntu 24.04 LTS server with a 32-core AMD EPYC-Milan CPU and 64 GB RAM.
+With the computation on this hardware and LLM calls to the GPT-3.5 API.
+In general, we use GPT-3.5 as the default LLM for our evaluation since it is the most used one by the selected approaches.
 6.4.2. Configuration and Code Changes
-For all parsers we set the initial temperature of the LLMs to 00 for a more deterministic output.
-Note that some parsers increase the temperature in case of a template mismatch or other reasons for some of their LLM calls.
 As mentioned in Sec.
-4.13.1, there are some issues with some of the parsers’ code which had to be resolved for the benchmark.
+5.13.1, there are some issues with some of the parsers’ code which had to be resolved for the benchmark.
 To ensure the comparability of the output of the parsers, other changes were also necessary to the code of each parser.
@@
 To ensure model or API compatibility, we updated deprecated versions of the OpenAI (python) package to the newer working ones and we added the functionality to call the OpenAI, Ollama, and TogetherAI API if not given.
-To have consistent output format, we ensure that the template placeholder symbols are always <*> and that the output is always a CSV file with at least the columns “Content”, and “EventTemplate”.
-“Content” refers to the log filtered by the log format (i.e.
-the log message).
-Since some parsers do not only parse the log content but the entire log, we modify the input to these parsers to only parse the content as well.
+To have consistent output format, we ensure that the template placeholder symbols are always <*>.
+Since the parsers LLM-TD (Vaarandi and Bahsi, 2024) and LogPrompt (Liu et al., 2024) do not only parse the log content but the entire log, we modify the input to these parsers to only parse the content as well.
 As found by He et al.
 (He et al., 2016b), this typically improves the parsing accuracy.
-Specific changes were made to DivLog (Xu et al., 2024) which is one of the earliest works on LLM-based log parsing.
+Specific changes were made to DivLog (Xu et al., 2024), which is one of the earliest works on LLM-based log parsing.
 It does not feature a caching mechanism and therefore, calls the LLM for each log line.
@@
 If a template cannot be matched with a log it is again sent to the LLM.
-We limit the maximum number of re-prompts to $3$ to avoid infinite LLM calls.
+We limit the maximum number of re-prompts to $3$ to avoid infinite LLM calls (which occurred in early experiments).
 LogPrompt also required some code corrections, other than typos.
@@
 Multiple works from our selection (Jiang et al., 2024a; Xu et al., 2024) as well as (Le and Zhang, 2023b) state that having a diverse candidate set is crucial to reduce the risk of overfitting to a specific log template, which is why they create specific sampling algorithms to sample from labeled logs.
-However, having a large proportion of templates at hand is an unrealistic setting in real-life application and requires users to label logs manually (Huang et al., 2024).
+However, this demands having a large proportion of templates at hand and may require users to label logs manually, thus may hinder real-life application (Huang et al., 2024).
 To ensure comparability, we exchanged the sampling processes of the supervised parsers DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) with a simple custom sampling approach: from all labeled logs of a dataset we randomly select $n$ (unique) templates.
@@
 3.
-For some reason, the implementation of ULP (Sedki et al., 2022) outputs the templates in a special format which is why the matching process of the evaluation function does not recognize correctly parsed logs, thus PA and FTA are 00 (GA is not affected).
+We can see that neither of the baseline parsers achieves satisfactory performance on FTA and PA, meaning that most templates are not matched exactly nor correct when compared token-by-token.
 Figure 3. Performance of the baseline parsers on the corrected LogHub dataset including Audit.
@@
 4(c).
-Note that we excluded LogPrompt from the evaluation with DeepSeek R1 because the LLM was barely able to extract any templates, resulting in scores of roughly 00 in experimental runs.
+Note that we excluded LogPrompt (Liu et al., 2024) from the evaluation with DeepSeek R1 because the LLM was barely able to extract any templates, resulting in scores of roughly 00 in experimental runs.
 The LLM did not provide the output in the structured format that was requested.
@@
 CodeLlama visibly performs worst, but it is also by far the smallest of the models with 7 billion parameters.
-Given the size difference of GPT-3.5 and DeepSeek R1, it is surprising that it does not outperform GPT-3.5, given its size.
+Given the size difference of GPT-3.5 and DeepSeek R1, it is surprising that DeepSeek R1 does not outperform GPT-3.5.
 Furthermore, we report the exact numbers for the evaluation with GPT-3.5 in Table 5 to show how the parsers performed on each individual dataset.
-Table 5. Performance of the selected parsers with GPT-3.5 on the corrected LogHub datasets including Audit
+Table 5. Performance of the selected parsers with GPT-3.5 on the corrected LogHub datasets including Audit. The best scores for each metric on each dataset are marked bold.
+Dataset
 Android
@@
 Baseline
-AEL GA 0.77 1.00 0.96 1.00 0.90 0.87 0.57 0.40 0.76 0.54 0.25 0.97 0.90 0.94 0.69 0.92 0.14 0.74
-FTA 0.44 0.50 0.22 0.53 0.46 0.24 0.10 0.27 0.17 0.23 0.08 0.44 0.28 0.20 0.27 0.45 0.00 0.29
+AEL GA 0.77 1.00 0.96 1.00 0.90 0.87 0.57 0.40 0.76 0.54 0.25 0.97 0.90 0.94 0.69 0.92 0.21 0.74
+FTA 0.46 0.50 0.22 0.53 0.46 0.24 0.10 0.27 0.17 0.23 0.08 0.44 0.28 0.20 0.27 0.45 0.00 0.29
 PA 0.39 0.69 0.34 0.62 0.68 0.51 0.16 0.17 0.17 0.25 0.02 0.67 0.38 0.04 0.15 0.75 0.00 0.35
-NED 0.90 0.92 0.84 0.93 0.97 0.70 0.56 0.74 0.71 0.92 0.60 0.97 0.71 0.71 0.90 0.92 0.38 0.79
+NED 0.90 0.92 0.84 0.93 0.97 0.70 0.56 0.74 0.71 0.92 0.60 0.97 0.71 0.71 0.90 0.92 0.49 0.79
 Brain GA 0.86 1.00 0.99 1.00 0.94 0.95 1.00 0.35 0.94 1.00 0.49 0.53 1.00 0.97 1.00 0.99 0.21 0.84
-FTA 0.37 0.67 0.24 0.80 0.46 0.20 0.40 0.40 0.33 0.23 0.21 0.74 0.44 0.34 0.45 0.57 0.00 0.40
+FTA 0.37 0.67 0.24 0.87 0.46 0.20 0.40 0.40 0.33 0.23 0.21 0.74 0.44 0.34 0.45 0.57 0.00 0.41
 PA 0.24 0.70 0.41 0.96 0.66 0.16 0.25 0.17 0.34 0.28 0.11 0.70 0.38 0.04 0.46 0.78 0.00 0.39
 NED 0.84 0.92 0.84 1.00 0.88 0.64 0.88 0.82 0.86 0.95 0.96 0.97 0.96 0.82 0.92 0.94 0.41 0.86
-Drain GA 0.83 1.00 0.96 1.00 0.89 0.95 0.78 0.42 0.79 0.79 0.22 0.76 0.92 0.96 1.00 0.97 0.14 0.79
-FTA 0.52 0.50 0.22 0.53 0.39 0.30 0.11 0.40 0.20 0.44 0.02 0.41 0.40 0.25 0.41 0.52 0.00 0.33
+Drain GA 0.83 1.00 0.96 1.00 0.89 0.95 0.78 0.42 0.79 0.79 0.22 0.76 0.92 0.96 1.00 0.97 0.21 0.79
+FTA 0.54 0.50 0.22 0.60 0.39 0.30 0.11 0.40 0.20 0.44 0.02 0.41 0.40 0.25 0.41 0.52 0.00 0.34
 PA 0.55 0.69 0.34 0.63 0.65 0.51 0.23 0.18 0.22 0.51 0.02 0.68 0.38 0.05 0.46 0.79 0.00 0.41
-NED 0.90 0.92 0.84 0.93 0.96 0.79 0.61 0.78 0.76 0.97 0.51 0.96 0.96 0.81 0.86 0.94 0.41 0.82
-SPELL GA 0.86 1.00 0.79 1.00 0.65 0.78 0.64 0.15 0.76 0.56 0.26 0.53 0.90 0.84 0.99 0.96 0.14 0.69
-FTA 0.25 0.50 0.06 0.43 0.39 0.15 0.10 0.14 0.05 0.23 0.00 0.04 0.19 0.15 0.10 0.37 0.00 0.18
+NED 0.90 0.92 0.84 0.93 0.96 0.79 0.61 0.78 0.76 0.97 0.51 0.96 0.96 0.81 0.86 0.94 0.43 0.82
+SPELL GA 0.86 1.00 0.79 1.00 0.65 0.78 0.64 0.15 0.76 0.56 0.26 0.53 0.90 0.84 0.99 0.96 0.34 0.71
+FTA 0.25 0.50 0.06 0.43 0.39 0.16 0.10 0.14 0.05 0.23 0.00 0.04 0.19 0.15 0.10 0.37 0.00 0.19
 PA 0.15 0.69 0.20 0.30 0.53 0.11 0.15 0.09 0.03 0.19 0.00 0.48 0.32 0.03 0.00 0.75 0.00 0.24
-NED 0.81 0.92 0.67 0.94 0.84 0.49 0.52 0.75 0.66 0.86 0.50 0.93 0.67 0.71 0.77 0.94 0.38 0.73
+NED 0.81 0.92 0.67 0.94 0.84 0.49 0.52 0.75 0.66 0.86 0.50 0.93 0.67 0.71 0.77 0.94 0.40 0.73
 ULP GA 0.74 1.00 0.93 1.00 0.95 0.99 0.90 0.18 0.81 0.43 0.47 0.02 0.92 0.68 0.41 0.99 0.34 0.69
-FTA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-PA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-NED 0.63 0.68 0.65 0.58 0.66 0.68 0.74 0.58 0.66 0.75 0.64 0.76 0.67 0.70 0.43 0.76 0.50 0.65
+FTA 0.29 0.00 0.30 0.00 0.53 0.20 0.33 0.40 0.28 0.07 0.04 0.38 0.20 0.44 0.38 0.38 0.00 0.25
+PA 0.23 0.00 0.38 0.00 0.85 0.16 0.07 0.20 0.24 0.15 0.02 0.51 0.02 0.07 0.39 0.50 0.00 0.22
+NED 0.67 0.74 0.69 0.61 0.77 0.72 0.80 0.61 0.69 0.78 0.66 0.78 0.77 0.74 0.44 0.82 0.51 0.69
 Unsupervised
@@
 The usefulness of reasoning models for log parsing should therefore be investigated in future research.
-Figure 5. Averaged performance of all parsers for CodeLlama, GPT-3.5 and DeepSeek R1.
+Figure 5. Averaged performance of LLM-based parsers for CodeLlama, GPT-3.5 and DeepSeek R1.
 7.1.4. Performance Difference on LogHub and corrected LogHub
@@
 An analysis of the performance of all parsers on the Audit dataset reveals that only the supervised parser, LILAC (Jiang et al., 2024a), and to some extent also DivLog (Xu et al., 2024), are able to parse the logs effectively.
-It should be noted that the configuration of the baseline parsers for Audit was not specifically tuned.
-It appears that the intended templates’ format of the logs is not straightforward for GPT-3.5 without supervised demonstrations.
+Despite the fact that the parameters of the baseline parsers were hyperparameter tuned, they did not achieve satisfactory results.
+It appears that the intended templates’ format of the logs is not straightforward for conventional parsers, but neither for GPT-3.5 without supervised demonstrations.
 This finding underscores the efficacy of supervised parsers, demonstrating that a mere two templates are sufficient to attain a reasonable performance level, potentially even a single one demonstrating the preferred template style.
-Nevertheless, it should be noted that this is merely a solitary example.
-Future research could establish a more extensive benchmark by incorporating a range of novel and diverse datasets.
-Figure 7. Performance of the selected parsers with GPT-3.5 only on the Audit data.
+Figure 7. Performance of the selected LLM-based parsers with GPT-3.5 and the baseline on the Audit data.
 7.2. Efficiency
@@
 As mentioned in Sec.
-4.13.1 and 5.4.2 there were some implementation problems with SelfLog (Pei et al., 2024).
+5.13.1 and 6.4.2 there were some implementation problems with SelfLog (Pei et al., 2024).
 They provided a faulty script for the efficiency evaluation which is why we ran the evaluation with the script they designed for the evaluation of the $2000$ log lines version of LogHub (Zhu et al., 2023).
@@
 RQ1: What are the main advantages and disadvantages of LLM-based log parsing approaches and non-LLM-based approaches?
-LLM-based log parsing approaches offer key advantages such as adaptability to diverse log formats, the ability to generalize across different system logs, and robustness to unseen log templates as demonstrated by the evaluation on the Audit dataset (Landauer et al., 2023b, 2021) and other work (Fu et al., 2022).
-Approaches like the unsupervised parser LogBatcher (Xiao et al., 2024) and the supervised parser LILAC (Jiang et al., 2024a) can handle complex and unstructured logs more effectively than traditional methods.
-However, they also come with drawbacks, including higher computational costs, potential hallucinations and overthinking for reasoning models (Cuadron et al., 2025), and difficulty in interpreting results due to their black-box nature.
-In contrast, non-LLM-based approaches, such as Brain (Yu et al., 2023) and ULP (Sedki et al., 2022), tend to be more interpretable, computationally efficient, and easier to deploy but lack the flexibility of LLMs in handling evolving log templates.
+LLM-based log parsing approaches offer key advantages such as adaptability to diverse log formats, the ability to generalize across different system logs, and robustness to unseen log templates as demonstrated by the evaluation on the Audit dataset (Landauer et al., 2023b, 2021) in Sec.
+7.1.5 and other work (Fu et al., 2022).
+Approaches like the unsupervised parser LogBatcher (Xiao et al., 2024) and the supervised parser LILAC (Jiang et al., 2024a) can handle complex and unstructured logs more effectively than conventional methods, as demonstrated in Sec.
+7.1.1 and Sec.
+7.1.2 and their corresponding publications (Jiang et al., 2024a; Xiao et al., 2024).
+They utilize both syntactic and semantic information within log data, outperforming methods that focus solely on syntactic (e.g.
+Drain (He et al., 2017), SPELL (Du and Li, 2016)) or semantic information (e.g.
+DivLog (Xu et al., 2024)).
+However, they also come with drawbacks.
+The usage of LLMs comes with a significant computational burden, implying the requirement of hardware capable of running LLMs locally or using external API services.
+Methods like caching can significantly reduce the number of LLM calls and therefore also the latency, but especially systems with a high number of templates or frequently changing templates (e.g.
+due to updates) may still use too many resources for certain use cases.
+The use of LLMs also introduces randomness into the process due to the randomness within the LLM’s output and possibly overthinking for reasoning models (Cuadron et al., 2025).
+In contrast, approaches that do not integrate language models into their process, such as Brain (Yu et al., 2023) and ULP (Sedki et al., 2022), have deterministic output, are computationally efficient, and are often easier to deploy but lack the flexibility of LLMs in handling diverse log templates.
 RQ2: To what extent do LLM-based log parsing approaches rely on labeled data and manual configuration?
 LLM-based log parsing approaches typically require labeled datasets for fine-tuning or evaluation, though some methods leverage self-evolutionary approaches using previously parsed templates to substitute the need for a priori labels.
-However, with only $2$ and $4$ shots of labeled logs, LILAC (Jiang et al., 2024a) achieves superior performance on the corrected LogHub dataset (Khan et al., 2022) and Audit (Landauer et al., 2023b, 2021), compared to the conventional parsers and other LLM-based parsers, excluding LogBatcher (Xiao et al., 2024).
+However, with only $2$ and $4$ shots of labeled logs, LILAC (Jiang et al., 2024a) achieves superior performance on the corrected LogHub dataset (Khan et al., 2022) and Audit (Landauer et al., 2023b, 2021), compared to the conventional parsers and other LLM-based parsers, excluding LogBatcher (Xiao et al., 2024) (Sec.
+7).
 Furthermore, LogBatcher demonstrates impressive performance on both runtime and effectiveness while being unsupervised and only requiring the log format as input parameter.
-Fine-tuning approaches also report outstanding performance, yet their stronger reliance on labeled logs and computational power and the robust and high performance of LogBatcher and LILAC demonstrate that ICL is a cheaper and performant alternative.
-In comparison to the baseline parsers AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023), the LLM-based parsers of our evaluation selection require on average less configuration parameters.
+Fine-tuning (Karanjai et al., 2024; Ma et al., 2024a; Mehrabi et al., 2024; Pang et al., 2024; Zhong et al., 2024) or even pretraining approaches (Ji et al., 2024) also report outstanding performance, yet their stronger reliance on labeled logs and computational power and the robust and high performance of LogBatcher and LILAC demonstrate that ICL is a cheaper but still performant alternative.
+Compared to the baseline parsers AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023), the LLM-based parsers of our evaluation selection require on average less configuration parameters (Sec.
+5.5).
 This finding is particularly noteworthy in light of the satisfactory performance achieved by LogBatcher (Xiao et al., 2024).
-This finding suggests that the deployment of LLMs can effectively substitute a substantial proportion of manual effort, thus enhance usability.
+This finding suggests that the deployment of LLMs can effectively substitute a substantial proportion of manual effort, and thus enhance the usability.
 RQ3: Which techniques can enhance the efficiency or effectiveness of LLM-based log parsing?
 It is evident that ICL and fine-tuning present powerful learning paradigms for LLM-based log parsing.
-They are not mutually exclusive to each other and can be employed simultaneously to improve parsing accuracy.
+They are not mutually exclusive to each other and can be employed simultaneously to improve parsing accuracy (Sec.
+5.4).
 ICL can be enhanced with RAG or with previously parsed or labeled logs as demonstrations.
-The ablative studies of reviewed papers (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) unanimously conclude that dynamic demonstrations and sophisticated demonstration selection algorithms improve the accuracy scores.
-Smart caching solutions have been shown to lead to efficiency improvements, in terms of runtime and monetary expenses, since LLM calls are usually costly.
-Template revision methods that correct, delete or merge related templates retrospectively can further improve cached templates.
-This can also help adapt the parser to evolving logs due to updates or other behavioral changes of the computer systems without reconfiguration or retraining.
-Evaluations show that is also possible to improve the effectiveness of parsers with more powerful LLMs, yet the effective application of reasoning models like DeepSeek R1 should be investigated more closely.
-It has been shown that variable aware prompts improve the effectiveness in direct parsing with LLMs.
-Naturally, more explanations in the prompt mean more tokens per parsed log which negatively affects the inference time of LLMs.
-Short and concise instructions and explanations could thereby alleviate this issue.
-Furthermore, creating coarse-grained clusters to capture the commonalities of logs and fine-grained clusters to capture the variabilities of logs to support the template extraction process constitutes another promising technique, that can be applied by conventional and LLM-based parsers alike.
-RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
-Key experiment design choices include the selection of benchmark datasets, evaluation metrics, training methodologies, and hardware specifications.
-Consistency in log datasets and pre-processing methods is crucial for fair comparisons.
-For instance, not all parsers require the log format as an input parameter, such as LLM-TD (Vaarandi and Bahsi, 2024) and LUNAR (Huang et al., 2024).
+The ablative studies of reviewed papers (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) unanimously conclude that dynamic demonstrations and sophisticated demonstration selection algorithms improve accuracy scores (Sec.
+5.6).
+Smart caching solutions show efficiency improvements in terms of runtime but also monetary expenses since LLM calls are usually costly (Sec.
+5.7).
+Template revision methods that correct, delete, or merge related templates in postprocessing can further improve cached templates.
+This can also help adapt the parser to evolving logs due to updates or other behavioral changes of the computer systems without reconfiguration or retraining (Sec.
+5.9).
+Evaluations show that it is also possible to improve the effectiveness of parsers with more powerful LLMs, yet the effective application of reasoning models like DeepSeek R1 should be investigated more closely (Sec.
+7).
+It has been shown that variable-aware prompts improve the effectiveness in direct parsing with LLMs.
+Naturally, more explanations in the prompt mean more tokens per parsed log, which negatively affects the inference time of LLMs.
+Short and concise instructions and explanations could thus alleviate this issue.
+Furthermore, creating coarse-grained clusters to capture the commonalities of logs and fine-grained clusters to capture the variabilities of logs to support the template extraction process constitutes another promising technique, that can be applied by conventional and LLM-based parsers alike (Sec.
+5.8).
+RQ4: Which experiment design choices hinder the comparability, and hence which guidelines should be adopted in terms of configuration, datasets used, evaluation metrics, and reporting to make results comparable?
+Consistency in preprocessing methods is crucial for fair comparisons.
+For instance, not all parsers require the log format (see definition in Sec.
+2) as an input parameter, such as LLM-TD (Vaarandi and Bahsi, 2024) and LUNAR (Huang et al., 2024).
 LUNAR explicitly includes log headers, such as IP addresses, timestamps, or levels, to grasp the full context of the log.
 For instance, partitioning logs into groups based on their timestamps is also possible with time intervals (Zhou et al., 2024).
-Previous research (He et al., 2016b) found that parsing only the extracted log message part, according to the log format, improves effectiveness, but especially LLMs with their generalization capabilities could utilize this extended context.
-A set of metrics such as GA, PA, FTA, but also NED (or ED), should be standardized across studies to cover all relevant characteristics of templates, since each metrics maps to specific requirements of specific downstream tasks.
-The results of our evaluation show high variance between these metrics for different parsers, indicating that the selection covers a variety of aspects.
-The choice of model hyperparameters also significantly impacts comparability.
-For example, the researchers of Lemur (Zhang et al., 2025) report near-perfect scores of $0.999$ and $0.996$ for FGA and GA on LogHub (Zhu et al., 2023), yet they require log format, regular expressions, and four other (hyperparameter-tuned) parameters for each dataset.
-This demonstrates the potential performance levels that can be achieved.
-However, the number of parameters required is impractical for real-world applications and a comparison to the performance of other parsers is not feasible.
-The aforementioned experiment design choices concern all types of parsers.
-Specific to LLM-based parsers, the employed LLM is also an important factor to consider regarding comparability of evaluation results.
-As demonstrated, CodeLlama, GPT-3.5, and DeepSeek R1 perform significantly different on the effectiveness scores.
+Previous research (He et al., 2016b) found that parsing only the extracted log message part, according to the log format, improves effectiveness, but especially LLMs with their generalization capabilities could utilize this extended context (Huang et al., 2024).
+In general, the choice of configuration parameters significantly impacts comparability.
+For example, the researchers of Lemur (Zhang et al., 2025) report near-perfect scores of $0.999$ and $0.996$ for FGA and GA on LogHub (Zhu et al., 2023), yet they require log format, regular expressions, and four other hyperparameter-tuned parameters for each dataset.
+This demonstrates the potential performance levels that can technically be achieved.
+However, the number of parameters required is impractical for real-world applications and complicates meaningful performance comparisons.
+The datasets related to LogHub (Zhu et al., 2023; Jiang et al., 2008; Khan et al., 2022) are used in most publications, namely $25$ out of $29$.
+LogHub and its versions are widely used benchmarks, enabling direct comparison with a large body of existing research and providing a standardized way to evaluate parsing performance.
+We therefore recommend using at least one of the LogHub versions for the sake of comparability.
+However, relying solely on LogHub datasets may not cover all possible log structures and scenarios.
+Incorporating custom or other open-source datasets reveals how well parsers generalize to different types of logs, identifies their strengths and weaknesses, and avoids overfitting to a single dataset (or dataset collection) (Sec.
+5.10).
+We recommend a set of metrics, namely GA, PA, FTA, and NED, to be standardized across studies to cover the relevant characteristics of templates, since each metrics maps to specific requirements of specific downstream tasks (Sec.
+5.11).
+This is important to prevent single-sided metric selection that focuses on single characteristics.
+Using FGA over GA should also be considered if class imbalance is a concern, but if not, there is little difference whether GA or FGA is used as long as the aspect of grouping performance is covered.
+The results of our evaluation in Sec.
+7 show high variance between these metrics for different parsers, indicating that this selection covers a variety of aspects.
+The aforementioned experiment design choices concern all types of log parsers.
+Specific to LLM-based parsers, the LLM employed is also an important factor to consider regarding the comparability of evaluation results.
+As demonstrated, CodeLlama, GPT-3.5, and DeepSeek R1 perform significantly different on the effectiveness scores (Sec.
+7).
 Additionally, variations in API versions, underlying model updates, and differences in temperature settings further contribute to comparability issues in performance across studies.
-This highlights the need for standardized benchmarking practices to ensure fair and reliable comparisons between different LLM-based parsing approaches.
 RQ5: To what extent are the presented LLM-based parsing methods accessible and reproducible?
-The accessibility and reproducibility of the presented LLM-based parsing methods are significantly hindered by issues related to code availability, documentation, and execution reliability.
+The accessibility and reproducibility of the presented LLM-based parsing methods are significantly hindered by issues related to code availability, documentation, and execution reliability (Sec.
+5.13.1).
 Many approaches do not provide public code, and even when code is available, it often lacks essential components, such as necessary scripts or instructions, making replication difficult.
@@
 Furthermore, discrepancies in the code, such as missing scripts for crucial steps or incorrect default settings, complicate reproducibility.
-Even when modifications are made to ensure compatibility — such as fixing typos, updating libraries, or implementing minor adjustments to improve consistency — these changes highlight the necessity of external intervention to make the methods functional.
+Even when modifications are made to ensure compatibility — such as fixing typos, updating libraries, or implementing minor adjustments to improve consistency (Sec.
+6.4.2) — these changes highlight the necessity of external intervention to make the methods functional.
 The need for such corrections aligns with broader findings from Trisovic et al.
@@
 further found that only a small proportion of the running codebases also produce the claimed results of the related papers (Olszewski et al., 2023).
-Additionally, some parsers require substantial configuration efforts, further limiting accessibility.
+Additionally, some parsers require substantial configuration efforts, further limiting accessibility (Sec.
+5.5).
 Whilst minor improvements, such as adding caching mechanisms to DivLog (Xu et al., 2024) or modifying LogPrompt’s (Liu et al., 2024) prompt size, enhance usability, they also introduce deviations from the original implementations, raising concerns about reproducibility.
-The discrepancy between our own evaluation results and the reported results raises further concerns about the validity of the publications’ evaluations but naturally also about our own evaluation.
+The discrepancy between our own evaluation results (Sec.
+7) and the reported results raises further concerns about the validity of the publications’ evaluations but naturally also about our own evaluation.
 Future benchmarks are expected to address these concerns.
-Finally, while log parsers based on LLM show great potential, their current state significantly hinders accessibility and reliable reproduction without substantial external effort.
+Finally, while log parsers based on LLM show great potential, their implementations’ current state significantly hinders accessibility and reliable reproduction without substantial external effort.
 8.2. Threats to Validity
@@
 Alternatively, it is possible that other factors, such as the utilization of different language models, are responsible for the observed variations.
-However, given that the alterations were predominantly minor, we anticipate that the correctness was not negatively influenced.
-The code has been made available on GitHub for replication.
+However, given that the alterations were predominantly minor, we anticipate that the correctness was not significantly negatively influenced.
+The code has been made available on GitHub for reproduction.
 Should any flaws in the adoption of the parsers be identified, we request to report an issue.
@@
 Therefore, we repeated the evaluation three times and computed the average of the measured values.
+8.2.3. Data Leakage
+Following previous work (Xiao et al., 2024; Jiang et al., 2024a) we identify data leakage through memorized logs from the pretraining phase of the LLMs as a threat to the validity of our evaluation results.
+However, (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024) report significant performance improvements with ICL compared to direct zero-shot parsing, suggesting a low likelihood of GPT-3.5-Turbo memorizing single log templates.
+In a small experiment we asked CodeLlama and DeepSeek R1 if they can output exact logs of the LogHub repository.
+In both cases they were not able to print exact logs, that matched any of those from the repository.
+The same held true when asked for templates.
+Diverse studies (Schwarzschild et al., 2024; Lu et al., 2024) suggest that direct fact memorization is unprobable and that the models are only likely to memorize often encountered text from diverse sources from their training (Schwarzschild et al., 2024).
 9. Conclusion
@@
 Sophisticated dynamic demonstration selection and caching strategies have been shown to significantly impact performance, reducing both LLM inference time and cost.
-However, despite these improvements, only two out of seven LLM-based parsers, namely LILAC (Jiang et al., 2024a) and LogBatcher (Xiao et al., 2024), are able to outperform the non-LLM-based parsers that constitute our baseline.
+However, despite these improvements, only two out of seven LLM-based parsers, namely LILAC (Jiang et al., 2024a) and LogBatcher (Xiao et al., 2024), clearly outperform the non-LLM-based parsers that constitute our baseline.
 Furthermore, we identified a lack of reproducibility and comparability, as many implementations lack publicly available code or datasets, comprehensive documentation, or consistency with reported results.
@@
 We hope that our systematic review, benchmark study, and insights contribute to the development of more effective, transparent, and user-friendly log parsing solutions.
-To facilitate future research and reproducibility, we make our evaluation results and source code publicly available on our GitHub repository.
+To facilitate future research and reproducibility, we make our evaluation results and source code publicly available on our GitHub repository (Beck, 2025).
 Acknowledgements.